A Complete Bibliography of Publications in the Journal of Mathematical Cryptology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 April 2024 Version 1.08

Title word cross-reference

(2, 2) [PS20]. (2, 2, 2) [AJZT19]. (n + 1) [ZMD20]. (p^{r} − 1)/3 [BW09]. 1 [SW07a]. 2 [AJZT19, Gau07, Hit09, SW07a]. 3 [CC15, LL15, MW12]. 4 [Kar10]. 5 [HL19]. 6 [Kar10]. _{LWE} [COP24]. ax ≡ b (mod n) [GP13]. GL_{2}(F_{p^{n}}) [TNS20]. k [AMW07, BW09]. F24m^{∗} [Kar10]. F36m^{∗} [Kar10]. Fp [BW09]. F_{p^{2}} [DM20]. Fpk

[Duq11]. Q(ζ2s) [BS19]. Z/NZ [ST24]. M [WP11]. Q [WP11]. MST_{3} [BCM09]. n [DSS20, ZMD20]. N = p^{r}q^{l} [LPS17]. p [KMNN19]. p + q [MA17]. ρ [Oka12]. MST_{3} [SvT10]. S [ZMD20].

-ary [KMNN19]. -bit [ZMD20]. -boxes [ZMD20]. -CP-ABE [COP24]. -error [AMW07, BW09]. -ideal [AJZT19]. -rotation [CC15]. -round [SW07a]. -server [SW07a]. -threshold [PS20]. -torsion [MW12]. -values [Oka12].

2019 [DKKS20]. 256 [DT17].

7 [GZ14].

ABE [COP24]. abelian [Rei17]. aborts [FS11]. Abu [Bla10]. Access [KK23, DS17, GD13, JSN13]. achieving [AL11]. action [JLLRL20, RS22]. adaptive [AL11, FS11]. add [FJ13]. add-rotate-xor [FJ13]. addition [FJ13]. additive [BDJ14]. admit [DN08]. Advanced [Sch08]. adversary [Dow15]. AE [ASB^{+}18]. Affine [CLS16, HWCD11]. AG [Tim24]. Against [Joy20, BS19, BBP^{+}20, Moc23, MSP12, WS09]. aggregation [HS15b]. Agrawal [PS08]. agreement [Bro23, Mul11]. AJPS [CG20]. al. [MU10]. algebra [BCSV20, KU18, RZ18]. Algebraic [BF09, Par18, RUS23, TKF^{+}20]. algebras [Ano20]. algorithm [BGMP08, BP20, BDJ14, JLLRL20, Pop17]. Algorithms [CCH^{+}20, AR15, DSG21, Eke21, FJ13, NV08, PP18]. alternating [SW15]. alternative [CCI24]. Among [KT20]. Analogue [DSGKS20b]. Analysis [KU15, KU18, Maz12, YYS^{+}17, CDK18, HVV14, MU08, NM08, Nan09, OPSB13, PSU13, Sch08, Wun19]. analyze [FJ13]. Annual [CLS20, CLY20]. anonymous [SSA17]. Aperiodic [BdW12, SvT13, vT18]. Application [DKKS20, Duq11, GGD20, Tim24, Tis11]. Applications [BBPS20, GPR17, BP20, FL08, FS11, JN16, SPB22]. approach [BFP09, CAR20, CLS16]. approaches [HS15b, TKF^{+}20]. Approximate [CCH^{+}20, Laa20]. approximations [SS16a]. argument [Mie08]. Ariffin [Bla10]. arithmetic [ATW08, Duq11, Gau07, HST10, HL09]. arrays [MN07]. ary [KMNN19]. associated [CN22]. asymmetric [AR15, Jia14]. asymptotic [KMPS10]. Attack [KMU20, TW20, AK17, BP22, BBP^{+}20, BF09, HF07, Mor08, MU10, Sch17, Wun19]. Attacks [Joy20, BS19, CW22, DHS20, GKL15, HL10, KMPS10, KHK10, Moc23, MSP12, MU08, PS08, Pou16, RUS23, RMS24, WS09]. attribute [AL11]. auction [KMNN19]. Authenticated [BLLN21, TSJL20, ASB^{+}18, CN22]. authentication [ACP10, Esk24, FHLMW08, JCK^{+}18, LU08, MU10, PS20]. automated [Kob07, Kob12]. Automatic [RMS24]. auxiliary [Kus18]. avalanche [LC07]. average [AL07].

Balanced [Rei21, CLS09]. barrier

[MPST16]. Based [PGS20, TSJL20, ACP10, BCM11, BGGJ20, CDK18, CG20, Esk24, GKL15, Gau07, GPV24, HS08, Jia14, KH23, KU15, KU18, LMPW15, MU08, PS15, Per12, RZ18, RST07, SS09, UJ20]. bases [BF09, FS08]. Beat [DM20]. Bent [Rei21, FM13]. Bergman [BT12]. Better [PP18]. between [BBP^{+}20]. bias [DS18]. Biggs [Bla09]. binary [ATW08, LMPW15, SW07a]. birthday [HF07, MPST16, MSP12]. Bit [TW20, ZMD20]. black [YH24]. black-box [YH24]. blind [FS11]. blinding [DSG21]. BLISS [TW20]. block [DR07, SS17, Sch08, ZLA21]. Boolean [BC12, CJST16, CLS09, CC15, CLS16, LZ23, ZLA21, ZHM^{+}22]. boomerang [RMS24]. Bound [DM20, Kus18]. bounds [BDJ14, SS17, YYTK20]. box [YH24]. boxes [ZMD20]. breaking [MPST16]. Brezing [Yoo15]. Bringer [CWZ12]. broadcast [AL07].

calculus [LL15, YYTK20]. Can [DM20]. Capitulation [AJZT19]. carry [FJ13]. carry-truncated [FJ13]. Cartesian [JCK^{+}18]. case [AK17, RST07]. CAST [DT17]. CAST-256 [DT17]. CBC [JN16]. CCA [KNP13]. CCA-secure [KNP13]. cells [Laa20]. certain [KU15]. Chabanne [CWZ12]. challenges [AK17]. Challenging [MSP12]. channel [Sch08]. characterisation [HS15a]. Characteristic [MR20, HST10]. characterization [BGMP08, FM13]. Characterizing [DHS20]. Cheng [PS08]. CHIMERA [BGGJ20]. choosing [Yoo15]. cipher [SS17, Tis11]. ciphers [DR07, FMS09, Moc23, Sch08]. ciphertext [CLSV24]. ciphertexts [AL11]. circuit [BBP^{+}20]. circulant [AN20]. class [AR15, CGK13]. Classes [AÖY15, AJZT19, WL13]. classical [BBP^{+}20]. cloud [PSU13]. CM [JLLRL20]. code [KH23]. code-based [KH23]. codes [BBB^{+}18, EHKN24, FHLMW08, JCK^{+}18, LdV19, PS20, Per12, SPB22, Tim24]. coding [FL08, PSU13]. coefficient [LZ23, ZHM^{+}22]. colexicographically [HM07]. Collision [BDJ14]. collisions [BS07]. coloring [ADPS14]. combinatorics [MN07]. combined [DSG21]. Combining [BGGJ20]. Common [CCH^{+}20, HL10]. communication [PS15]. commutative [JCK^{+}18]. Compact [CQS11, Per12]. comparative [CDK18]. comparison [BK09]. complete [BGMP08, Oka12]. Complex [BCIV20]. Complex-Valued [BCIV20]. complexities [SS17]. Complexity [YYTK20, AMW07, BW09, DS17]. composite [CDK18]. compositeness [PR20]. Compression [BG16, Kar10]. computation [ADPS14, CS21, CI14, DGG^{+}15, Duq11, DEF14, GÓS07]. computational [Ano20]. computations [CDN18]. Computing [Bis11, Eke21]. concrete [APS15]. concurrent [WS09]. concurrent-reset [WS09]. condition [DSS20]. conferences [JP20]. confusion [LZ23, ZHM^{+}22]. conjecture [CLS09]. conjugacy [LU08]. conjugation [MU14]. connected [DS17]. connection [Zaj13]. Constant [TW20, AL11]. constant-size [AL11]. constants [Grz20]. constructed [ZMD20]. Constructing [CJS14]. Construction [vT18, AGH17, ADPS14, CL09, FHLMW08, IK23, JCK^{+}18]. constructions [GPR17, LdV19, PS15, WC07]. Converting [CDK18]. convolution [OS14]. Coprime [GP13]. correlation [DR07, DT17]. corresponding [BFJT12]. Cosets [KT20]. Coskun [AK17]. Counting [DN08, WL13]. CP [COP24, COP24]. CP-ABE [COP24]. criterion [LC07]. critical [Bla09]. CRT [CCH^{+}20]. CRT-variant [CCH^{+}20]. Cryptanalysing [Bla09, Mul11, Bla10]. Cryptanalysis [BCM09, BCM11, BKL22, CBW07, Esk24, LU08, LPS17, MU14, TL15, BP20, CG20, DT17, LNR09, MU10, PGS20, RST07, SS16a, SS17, SvT10, Tis11]. Cryptographic [BBPS20, CJST16, BCSV20, BFJN20, MP08, SPB22, SS09, ZMD20]. cryptography [CAR20, COP24, GGD20, GPR17]. Cryptol. [GZ14]. cryptology [CW22, Gut09, JP20]. Cryptosystem [PGS20, BCM09, Bla10, CBW07, CG20, HF07, Jus14, Jus15, KU15, SSS11, SvT10]. cryptosystems [AÖY15, CDK18, DJP14, EOS07]. CSIDH [BBP^{+}20]. cubic [CC15]. cumulative [MN07]. Curve [HKP^{+}20, KT20, CJS14, CQS11, DJP14, FK18, Tim24]. Curves [BGK^{+}20, BG16, Bis11, Box12, BDFM21, CDS20, DN08, DEF14, FSV09, GPRS09, GZ13, GZ14, HST10, HWCD11, Hit09, LS07, MW12, Mor08, Oka12, ST24, Sch17, Sha14, Yoo15]. cyclic [AJZT19]. cyclotomic [AMW07, DSS20, Kar10, Sha14].

DAGS [BBB^{+}18]. Data [BCIV20, CN22, SS17, SU14]. decentralized [Pop17]. decreasing [YYS^{+}17]. decryption [YYS^{+}14]. DeepBKZ [Yas20]. defeat [DSG21]. degree [Hit09, KW19]. degrees [TKF^{+}20]. Delegating [DKKS20]. delegation [ZSN20]. Deligne [Tim24]. Dembowski [AÖY15]. density [FS09]. Dependent [FGG^{+}20]. descent [KMPS10]. design [GPR17]. designated [YH24]. Designing [BBPS20]. detailed [Wun19]. determined [KK23]. Devastating [TW20]. Differential [CCI24, DT17]. differential-zero [DT17]. differentials [DR07]. Diffie [KM08, Kus18, Par18]. digital [RUS23]. dihedral [CS24]. dimension [EHKN24].

Dimensions [MR20]. Discrete [HKP^{+}20, MR20, BT12, Bla09, Bla10, BDJ14, CI14, Eke21, Gal12, KM08, Sho10]. Discretisation [MP20]. disguised [Mor08]. distinguisher [Gal12]. Distortion [GPRS09]. distributed [DGG^{+}15]. Distribution [LS07, BS07, Jus14, Jus15, Maz12, PS08, vzGS09]. Distributions [MP20, DR07]. Divisor [CCH^{+}20]. dragon [SSS11]. DSA [Pou16]. dual [NP19, Yas20]. duality [Csi20]. Dyadic [BBPS20, BBB^{+}18, Per12].

ECC [Joy20]. ECDLP [DM20, YYTK20]. Editor [CLS20]. Edwards [BG16]. effect [Mur12]. effectiveness [KMPS10, Mur12]. Efficiency [GPV24]. Efficient [BBPS20, DEF14, HST10, DGG^{+}15, FM13, SSS11, SW07b, WS09]. Efficiently [BCIV20, Bla09]. elementary [Kar20a, Kar20b, ZLA21]. elements [Yoo15]. Eleventh [JLNN20]. ElGamal [PGS20]. Elliptic [BGK^{+}20, HKP^{+}20, KT20, Bis11, Box12, CJS14, CQS11, DJP14, DEF14, FSV09, FK18, HST10, HWCD11, LS07, MW12, Mor08, Oka12, ST24, Sch17, Sha14, Yoo15]. EMAC [JN16]. embedding [Hit09]. encapsulation [BBB^{+}18]. encodings [NP19]. Encryption [BLLN21, BCIV20, BGGJ20, DS20, ASB^{+}18, AL07, AL11, CLSV24, CN22, CEM15, IK23, KNP13, NP16, SPSS12, TL15, YYS^{+}14]. endomorphism [Bis11]. Enhancing [SPS19]. EPIR [CWZ12]. equation [MP08, Zaj13]. equations [GP09]. Equidistribution [KT20]. equivalence [CLS16, PS20, WL13]. Equivalent [WP11]. Erratum [GZ14]. error [AMW07, BW09]. errors [APS15, BBGS19]. Establishment [TSJL20, BCM11]. Estimation [BBGS19]. Evaluation [DS20, CDGM14, CWZ12, TJBY13]. even [Kar20a, Kar20b]. Evolution [CW22, BGMP08]. Exact [HL19, YYS^{+}14]. examples [DC14]. Exchange [BGK^{+}20, RM20, Ala17, Jia14, KU18, Par18, RS22]. Exclusion [Rei21]. exhaustive [RZ18]. existence [WC07]. expansions [FS08]. experiments [CCI24]. explicit [ATW08, Grz20]. exploration [HWCD11]. exponent [HL10]. exponentiation [DSG21]. Exponentiations [DKKS20]. exponents [Böc09, Kar20a, Kar20b]. Exposure [DSGKS20a]. expressions [KMNN19]. Extension [Joy20]. extracting [SU14].

Factor [Kar10]. Factor- [Kar10].

Factoring [Sic20, PR20]. factorization [Zaj13]. fall [KW19]. Families [Hit09, MW12, Ano20, CL09, FK18, Oka12, SPB22, Sha14, WC07, Yoo15]. Family [BLLN21, ACP10, BFJT12, FM13]. Fast [Gau07, OS14, Duq11, HL09, HL10, MPST16, jWW12, vzGS09]. Fault [Joy20]. FCSRs [FMS09]. Fermat [Kar20b, Kar20a]. field [ATW08, HSWZ20, Kus18]. fields [ATW08, AJZT19, BFP09, GP09, HST10, LPS20, LC07]. filtered [OPSB13]. find [Rei21]. Finding [Gal12, BS19, PP18, Yas20]. finite [BFP09, CBW07, FL08, HST10, HSWZ20, JCK^{+}18, Kus18, LC07, Sho10]. first [KW19, CLY20]. Flattening [DS20]. flaw [ASB^{+}18]. Flips [TW20]. flow [SW07b]. Foreword [Gut09, Mag13]. forms [HS08]. formulæ [ATW08]. formulas [MA17]. foundation [PSU13]. framework [BP20, BCSV20]. Free [DS20, Ano20]. friendly [Box12, FK18, Oka12, Sha14, Yoo15]. Frontmatter [Ano17, Ano18d, Ano18a, Ano18b, Ano18c, Ano19c, Ano19a, Ano19b]. Fully [BGGJ20, Dow15, TL15, YYS^{+}14]. function [ACP10, GP09, Nan09, NSW09, PS15]. Functional [AL11]. Functions [Rei21, TNS20, BC12, CJST16, CDS20, CLS09, CC15, CLS16, FM13, Gau07, KMNN19, KHK10, LZ23, MSP12, OS14, SW15, SW07a, ZSN20, ZLA21, ZHM^{+}22]. fundamental [ASB^{+}18]. Further [LZ23].

general [Eke21, PSU13]. generalised [Dow15, MN07]. Generalization [DS18, Par18, Tim24]. generalized [KHK10]. generate [SW15]. Generating [FK18, vzGS13]. generation [ABD^{+}13, BGMP08, Pop17, vzGS09]. generator [BS19]. Generic [LdV19, Moc23]. Gentry [YYS^{+}14]. genus [ATW08, CDS20, GPRS09, Gau07, Hit09, LL15]. genus-2 [CDS20]. geometric [MP08]. geometries [JCK^{+}18]. geometry [FL08]. girth [HL19]. global [GP09]. Goldreich [SPS19]. Goldwasser [Jus14, Jus15, SPS19]. GPST [BP22]. Gram [YYS^{+}17]. graph [ADPS14, DS17, GD13]. graphs [CK20, HL19, JN16, PS15, PP18, Sho10]. Greatest [CCH^{+}20]. GRH [Bis11]. Gröbner [BF09]. Group [DKKS20, PGS20, ST24, Bla09, CBW07, CW22, HWCD11, JLLRL20, KU15, Kus18, RST07, SW15, Zaj13]. group-theoretic [CW22]. groups [GKL15, Rei17, SS09]. Grover [BP20]. GS [BBB^{+}18].

Halevi [SPS19, YYS^{+}14]. hard [HS08]. hardness [APS15, BS19, BBGS19]. Hash [CDS20, DSGKS20b, NSW09, TNS20, ACP10, AN20, CL09, FJ13, KHK10, MPST16, MSP12, OS14, WC07]. hash-function [ACP10]. hashing [FSV09, LMPW15]. Hellman [KM08, Kus18, Par18]. Herley [AK17]. Hessian [BDFM21]. Heuristics [Box12, CW22]. HFE [BFJT12]. hidden [CS24, SS09]. hidden-order [SS09]. higher [CS21]. Hints [Sic20]. HMAC [KM13]. HMQV [Men07]. Homomorphic [BCIV20, BGGJ20, DS20, CBW07, TL15, YYS^{+}14]. hull [Mur12]. Hybrid [BFP09, KH23, Wun19]. hyper [CW22, FM13]. hyper-bent [FM13]. hyper-heuristics [CW22]. hyperelliptic [ATW08, DN08, GZ13, GZ14].

ID [PS15]. ID-based [PS15]. Ideal [LPS20, AJZT19, BBP19, BS19, HS15a, JSN13]. idealness [Sha14]. Identification [HS08, AK17, SW07b, WS09]. II [Jus15, Kob12]. implementation [CDGM14, YYS^{+}14]. implications [FMS09, GÓS07]. impossibility [YH24]. improve [MA17]. Improved [CG20, Kus18, NM08, Nan09, PGS20]. Inclusion [Rei21]. Inclusion/Exclusion [Rei21]. increased [MSP12]. indefinite [HS08]. index [LL15, YYTK20]. Indifferentiability [MPST16]. Indirect [ABD^{+}13]. inequality [Grz20]. Infinite [DC14]. information [HL19, SW07a]. injection [ABD^{+}13]. inner [AL11]. input [DGG^{+}15]. insecure [Bro23]. insider [BBP19]. instantiations [LdV19]. Integer [PR20, HM07]. integers [BK09]. Interactive [BGK^{+}20, CEM15]. interplay [ATW08]. intersecting [SPS19]. intersection [Esk24]. Intervals [KT20]. inverting [JLLRL20]. Isogenies [BGK^{+}20, BDFM21, CDS20, CJS14, DJP14]. Isogeny [TSJL20, CK20, TKF^{+}20]. Isogeny-Based [TSJL20]. Isolated [Sch17]. isomorphism [HSWZ20]. Issue [DKKS20]. iterated [KHK10].

J [GZ14]. Jacobi [DEF14]. Jacobian [ATW08, Sho10]. JMC [DKKS20]. joint [NP16].

KASUMI [SW15]. Kayawood [KMU20]. Key [BBB^{+}18, BGK^{+}20, DSGKS20a, DS20, RM20, TSJL20, Ala17, BCM09, BCM11, BFJT12, Bro23, CDN18, CAR20, CEM15, CBW07, DS18, GGD20, Jia14, KU18, KNP13, Maz12, Mul11, Par18, PS08, PS15, RS22, SPSS12, SSS11, SvT10, TL15, WP11]. key-recovery [BFJT12]. keys [BFJT12, KMU20, Per12, WP11]. keystream [DS18]. keyword [CLSV24]. knowledge [SW07b]. known [ZMD20]. Koblitz [DN08].

large [Grz20]. Last [Kar20a, Kar20b]. Lattice [MR20, Pou16, Wun19, YYS^{+}17, Yas20]. lattice-reduction [Wun19]. lattices [Laa20, SPS19]. laws [HWCD11]. Leakage [CDGM14, FGG^{+}20, Ala17, CAR20, KNP13]. leakage-resilient [Ala17, CAR20, KNP13]. learning [APS15, BBGS19]. least [HL19]. Leftover [DSGKS20b]. Lemma [DSGKS20b]. Length [GKL15, RST07, MU08]. Length-based [GKL15, RST07, MU08]. lengths [YYS^{+}17]. Lightweight [BLLN21, GPR17]. Linear [HS15b, AMW07, BCSV20, BW09, CDN18, DT17, EHKN24, Mur12, RZ18]. linearly [OPSB13]. Log [MR20, KM08]. Logarithm [HKP^{+}20, BT12, Bla09, Bla10, Sho10]. logarithmic [BdW12, Rei17, SvT13, vT18]. logarithms [BDJ14, CI14, Eke21, Gal12]. look [HL09, Kob07, KM08, Kob12, KM13, Men07, SS16a]. low [ATW08]. lower [Kus18]. LPS [PP18]. Lusztig [Tim24]. LWE [BGGJ20, DSGKS20a, MP20].

MAC [ABD^{+}13, JN16]. MAKE [BKL22, RS22]. maps [GPRS09]. masking [Sch08]. Masthead [Ano13]. Math. [GZ14]. MathCrypt [CLS20, CLY20]. Mathematical [Gut09]. Matrices [PGS20, GPR17]. matrix [BCM11, DSS20, KU18, MU14, RS22, ZSN20]. matrix-based [BCM11]. matroids [MFP10]. maximum [EHKN24]. McEliece [EOS07, Per12]. McEliece-type [EOS07]. MDS [GPR17]. Mean [MU10]. Mean-set [MU10]. Medium [MR20]. meet [Wun19]. meet-in-the-middle [Wun19]. memory [LL15]. mental [jWW12]. Mersenne [CG20]. message [ABD^{+}13, Dow15, FHLMW08]. Method [Joy20, WL13, YYTK20, Yoo15, Zaj13]. Methods [JP20, DSG21, HL09, JCK^{+}18, Sch08]. mF [CN22]. Micali [Jus14, Jus15]. MICKEY [Tis11]. middle [Wun19]. Miller [Böc09]. Minimal [FS08, HM07]. mixing [LMPW15]. mnemonic [AGH17]. mode [ASB^{+}18, CN22]. model [BBP19, DN08]. modified [HVV14]. modular [HL09]. moduli [LPS17]. modulo [Bla10]. modulus [HL10]. monodromy [Sho10]. Monomial [Rei21, CLS16]. monotone [CJST16]. MOV [Sch17]. MQQs [CGK13]. mR [COP24]. MR3101014 [GZ14]. MRHS [RZ18, Zaj13]. Multi [PSU18, FHLMW08, Hin08, YH24]. multi-designated [YH24]. multi-message [FHLMW08]. multi-prime [Hin08]. Multi-prover [PSU18]. multi-receiver [FHLMW08]. multicasting [Maz12]. Multicollision [KHK10]. multicubic [LPS20]. multilevel [JSN13]. Multiparty [BGK^{+}20, ADPS14, CS21]. Multiple [DT17, FM13]. multiplicative [Kus18]. multiplier [EHKN24]. multivariate [AÖY15, BFP09, CGK13, HF07, SSS11, TJBY13].

naive [YYTK20]. negation [AL11]. network [FL08, PS08]. networks [HS15b]. NIKE [UJ20]. NLFSRs [OPSB13]. Non [BGK^{+}20, TW20, BT12, CDN18, CEM15, Jus14, Jus15, KM08, Sha14]. Non-Constant [TW20]. non-idealness [Sha14]. Non-Interactive [BGK^{+}20, CEM15]. non-linear [CDN18]. non-representable [BT12]. non-residues [Jus14, Jus15]. non-standard [KM08]. Nonsmooth [Tis11]. norm [GP09]. normal [SS16a]. note [CS21]. notions [SPSS12]. NP [HS08]. NP-hard [HS08]. NTRU [BF09, DHS20, DS20]. number [AJZT19, BBGS19, BFJN20, DSS20, MA17, Pop17, WL13, JP20]. number-theoretic [BFJN20, JP20]. numbers [vzGS09]. Numerical [LNR09]. NutMiC [DKKS20].

Oblivious [TJBY13, FS11]. odd [HST10]. off [BBP^{+}20]. offs [LL15]. OMAC [Nan09]. One [BK09, TW20, Bla10, PS15]. One-round [BK09]. one-way-function [PS15]. opening [CEM15]. Operations [BBPS20, CCI24]. Optimal [PS15, Sil07, CDGM14]. optimality [CDN18]. oracle [BP20]. orbit [Gal12]. order [CDK18, SS09]. orders [Eke21]. Oribatida [BLLN21]. Orienting [CK20]. Ostrom [AÖY15]. overhead [AL07, BP20]. overstretched [DHS20].

pair [NP19]. pairing [Box12, CDK18, Duq11, FK18, GÓS07, Oka12, Sha14, Sho10, Yoo15]. pairing-based [CDK18]. pairing-friendly [Box12, FK18, Oka12, Sha14, Yoo15]. pairings [DEF14, GZ13, GZ14]. parallel [CCI24]. parameterization [Sil07]. parameters [FK18]. Partial [DSGKS20a]. participants [DS17, GD13]. Password [TSJL20, Jia14]. Password-Authenticated [TSJL20]. password-based [Jia14]. path [PP18]. path-finding [PP18]. Perfect [GD13, CL09, WC07]. period [BW09]. permutation [BBP19, LMPW15]. permutations [AGH17]. Persistent [Jia14]. pipe [MPST16]. Pisot [FS08]. PKA [AR15]. Plactic [Bro23]. PMAC [NM08]. Points [KT20, BG16, LS07]. poker [jWW12]. Pollard [BDJ14]. Poly [SSS11]. Poly-dragon [SSS11]. Pólya [CLS16]. polycyclic [GKL15, KU15]. polycyclic-group-based [KU15]. Polylogarithmic [Mie08]. polymatroids [KK23, MFP10]. Polynomial [KMNN19, CWZ12, DSS20, JLLRL20, WL13, ZSN20]. Polynomials [HKP^{+}20, AÖY15, KW19, TJBY13]. PoRs [LdV19]. possessing [SU14]. post [COP24]. post-quantum [COP24]. Power [JLNN20, ACP10, TKF^{+}20]. practical [BFJT12, CAR20, NV08]. practice [HL10]. Predicate [NP19]. predistribution [PS15]. Preface [CLS20, CLY20, JP20]. Prefix [FGG^{+}20]. Prefix-Dependent [FGG^{+}20]. presence [Dow15, Sch08]. preserving [Esk24, ZSN20]. prime [CDK18, Hin08, TKF^{+}20]. primes [ACP10, FS09, vzGS13]. primitive [SS09, Yoo15]. primitives [BFJN20, CDN18]. Principal [LPS20]. Privacy [ZSN20, Esk24]. Privacy-preserving [ZSN20, Esk24]. private [DGG^{+}15, Esk24, HL10, KMU20, SW07a]. Probability [DR07]. Problem [CCH^{+}20, HKP^{+}20, LPS20, BT12, BBGS19, Bla09, Bla10, CS24, HSWZ20, Kus18, NV08]. problems [BCSV20, HS08, KM08, TKF^{+}20]. Proceedings [CLS20, CLY20]. processes [SW07a]. Processing [BCIV20]. Product [DKKS20, MP20, OS14]. products [AL11, IK23]. proof [Kar20a, Kar20b, NP19, PSU13, PSU18]. proof-of-retrievability [PSU13]. properties [CJST16, ZMD20]. property [EHKN24]. Protecting [Joy20]. protocol [CWZ12, KU18, KMU20, LU08, MU10, SW07b, jWW12, WS09]. protocols [AK17, BCM11, IK23, SW07a]. Provable [Moc23]. prover [PSU18]. proving [Kob07, Kob12]. proxy [SSA17]. Pseudo [Ano20, BGMP08, Pop17]. Pseudo-free [Ano20]. pseudo-random [Pop17]. pseudorandom [Nan09, vzGS09]. Public [SvT10, AL11, BCM09, CAR20, CBW07, GGD20, KNP13, SPSS12, SSS11, WP11]. public-attribute [AL11]. public-key [CAR20, CBW07, GGD20, SPSS12].

quadratic [AÖY15, CGK13, HF07, HS08, Jus14, Jus15]. Quantum [CI14, Eke21, BS19, BP20, BBP^{+}20, CJS14, COP24, DJP14, JLLRL20, RUS23]. quantum-resistant [DJP14]. quartic [AJZT19, DEF14]. Quasi [HKP^{+}20, Per12]. quasi-dyadic [Per12]. Quasi-subfield [HKP^{+}20]. quasigroups [CGK13]. query [SW07a]. quotient [MU08].

Rabin [Böc09]. Ramanujan [PP18]. Random [MU08, BGMP08, EHKN24, Pop17]. randomized [Böc09, CLSV24]. range [YYS^{+}14]. rational [MW12]. ratios [HL19]. RC4 [BGMP08, DS18]. realization [SvT10]. realizing [PS15]. receiver [FHLMW08]. reconstruction [SW07a]. Recovering [FGG^{+}20]. recovery [BFJT12]. Recursive [BC12, CL09]. reduced [DT17]. reduced-round [DT17]. reducing [BP20]. reduction [Wun19]. refinement [LS18]. regular [DSG21, MSP12]. relations [DS18, SPSS12]. relying [BS19]. Remarks [RM20, FMS09]. repairable [LS18]. representable [BT12]. Representation [DM20]. representations [CQS11, HM07]. Reproducible [SPB22]. requirements [NSW09]. research [LZ23]. reset [WS09]. Residue [JLNN20, CS21]. residues [Jus14, Jus15]. resilient [Ala17, CAR20, HS15b, KNP13]. resistance [MSP12]. resistant [DJP14]. restricted [BBGS19]. results [DS18, LZ23, SW07a]. Rethinking [ATW08]. Retraction [Kar20b]. retrievability [PSU13, PSU18]. retrieval [SW07a]. Revisited [Joy20, AN20, Laa20, NP16, SS16b]. Revisiting [JN16]. revocable [COP24]. rho [BDJ14]. Richelot [CDS20]. Rigorous [SS17]. Rijndael [RMS24]. Ring [DSGKS20b, Joy20, BT12, YH24, BGGJ20, DSGKS20a, MP20]. Ring-LWE [DSGKS20a, MP20]. Ring-LWE-based [BGGJ20]. Rings [PGS20, Bis11, JCK^{+}18]. Rivest [HVV14]. RNS [Duq11]. robust [JSN13, PS20]. Roos [DS18]. Root [DM20]. rotate [FJ13]. Rotation [Rei21, CC15, CLS16]. round [BK09, DT17, Mie08, SW15, SW07a]. RSA [DSG21, Hin08, HL10, LPS17, MA17].

safe [AK17, vzGS13]. same [SU14]. samples [BBGS19]. SAP [GGD20]. scheme [Esk24, HVV14, HSWZ20, KH23, Maz12, Mul11, PS08, SSA17, SPS19, TL15]. Schemes [BLLN21, BGGJ20, DKKS20, BBP19, BS19, GD13, HS15a, JSN13, LS18, MFP10, MU14, PSU13, PS15, PS20, Pou16, RUS23, SS16b, YH24]. Schmidt [YYS^{+}17]. Schnorr [NSW09]. search [CLSV24, MA17, RZ18, RMS24]. Searchable [CLSV24]. Second [CLS20, Gut09]. Secret [Csi20, DC14, GD13, HS15a, HL19, JSN13, MFP10, CEM15]. Secrets [FGG^{+}20]. Secure [Dow15, BK09, CS21, FL08, JSN13, KNP13, Maz12, PSU13, SSA17, SW07b, SW07a, SS16b, WS09]. Security [FS11, HVV14, OPSB13, ACP10, AL11, BBP19, CDGM14, EOS07, GÓS07, Hin08, Moc23, MPST16, NM08, Nan09, NP16, SPSS12, DSGKS20a]. Self [GZ13, Yas20, GZ14]. Self-dual [Yas20]. Self-pairings [GZ13, GZ14]. Semaev [YYTK20]. semidirect [IK23]. semigroups [CI14]. Sensitivities [ZLA21]. sensor [HS15b, PS08]. sequences [AMW07, BW09, LS07]. server [SW07a]. set [Esk24, Gal12, MU10]. setting [CDK18]. seven [DS17]. sharing [Csi20, DC14, GD13, HS15a, HL19, JSN13, MFP10]. shifted [LU08]. Short [LPS20, BS19, YYS^{+}17, Yas20]. shortest [NV08]. Sibert [MU10]. side [Sch08].

Sidelnikov [BW09]. SIDH [GPV24, UJ20]. SIDH-based [GPV24, UJ20]. Sieve [NV08, Grz20]. Sieving [MR20]. Sign [TW20]. Signature [DKKS20, HSWZ20, NP16, RUS23, SS16b, YH24]. signatures [BdW12, FS11, GPV24, HS08, NP19, NSW09, Rei17, SvT13, vT18]. Signcryption [BBP19, KH23, SSA17]. SIKE [BP20]. Simplified [GÓS07]. six [GD13]. size [AL11, BBP^{+}20]. SKENO [CEM15]. small [HST10, HL19, HL10, Hit09]. SNFS [Sil07]. solutions [GP13]. solve [Zaj13]. solver [RZ18]. solvers [LNR09]. solving [BFP09, Bla09, BDJ14, GP09, MP08, TKF^{+}20]. Some [FL08, FMS09, SW07a, AJZT19, BC12, DS18, FS09, HL09, HL10, IK23, LS07]. space [JLLRL20]. sparse [AK17, FK18]. Special [DKKS20, CGK13, FS09]. specified [PS15]. Square [DM20]. squared [YYS^{+}17]. squared-sum [YYS^{+}17]. squeezing [CDGM14]. Srivastava [Per12]. standard [KM08]. Statistical [BS07]. Stickel [Mul11]. Stochastic [DSG21, Sch08]. storage [PSU13]. stream [FMS09, Moc23, Tis11]. streams [DGG^{+}15]. Strict [LC07]. Strongly [SvT13, AR15, vT18]. structure [DS17, JN16, JSN13, ST24]. structures [GD13, KK23]. subexponential [ADPS14, CJS14, JLLRL20]. subexponential-time [JLLRL20]. subfield [HST10, HKP^{+}20]. subgroup [CS24]. subgroups [Kar10, MU08]. Submission [DKKS20]. subsequences [BW09]. Subset [vzGS09]. sum [HF07, YYS^{+}17, vzGS09]. summary [EOS07]. summation [KW19]. supersingular [BP22, CK20, DJP14, GPRS09]. superspecial [CDS20]. support [AL11]. Survey [GGD20, LS18]. Suzuki [Tim24]. Symbol [JLNN20]. symbols [CS21]. Symmetric [Rei21, CDN18, CLS09, CC15, CLS16, TL15, ZLA21]. System [RM20, NP19]. systems [BFP09, Mie08, WP11, Zaj13].

Takes [TW20]. Tame [Rei17]. technique [NP19]. Techniques [UJ20]. terms [FM13]. test [Böc09]. th [DSS20]. their [EOS07]. Theorem [Kar20b, CLS16, Kob07, Kob12, Kar20a]. theorem-proving [Kob07, Kob12]. theoretic [BFJN20, CW22, JP20]. Theory [CC15, GPR17, PSU13]. theta [Gau07]. Thompson [RST07]. Three [MR20, BCM11]. threshold [JSN13, LS18, PS20]. Tillich [TNS20]. Time [LL15, TW20, CJS14, JLLRL20]. Time-memory [LL15]. Timing [TW20]. torsion [MW12]. trace [BG16, EHKN24, FM13]. trade [BBP^{+}20, LL15]. trade-off [BBP^{+}20]. trade-offs [LL15]. tradeoffs [Eke21]. transfer [FS11]. transmission [AL07, Dow15]. Tropical [RM20, KU18]. truncated [FJ13]. trustless [Pop17]. twisted [BG16, BDFM21]. Two [LMPW15, PS08, GPRS09, Mie08, RUS23, SW07b]. two-flow [SW07b]. Two-permutation-based [LMPW15]. two-round [Mie08]. TWOOA [FHLMW08]. Type [TNS20, AJZT19, EOS07, Jus14, Jus15, Tim24].

uadratic [WP11]. ultivariate [WP11]. unbounded [DGG^{+}15]. uncloaking [KMU20]. Unconditionally [JSN13, SS16b, PSU13, SW07a]. Unconditionally-secure [JSN13]. uniform [KK23]. universal [ACP10, Ano20]. upper [SS17]. use [BP20]. Using [FJ13, Rei21, BBB^{+}18, BF09, CCI24, CDS20, CW22, FK18, Kus18].

v1.3 [BLLN21]. Valued [BCIV20]. values [Oka12]. Vandermonde [DSS20]. variant [CCH^{+}20, Grz20, LPS17]. variants [HL10, Mul11]. vector [NV08]. vectors [BF09, EHKN24, YYS^{+}17, Yas20]. verifiable [ZSN20]. verifiers [YH24]. via [CS21, DM20, NP19]. view [MP08]. Volume [CLS20, CLY20]. Voronoi [Laa20]. VSH [BS07].

way [PS15]. weak [AÖY15, BFJT12]. weight [FS08, HM07]. weighted [HS15a]. weights [BC12]. Weil [HF07, KMPS10]. Weng [Yoo15]. wide [MPST16]. witnesses [PR20]. Witt [BF09]. Workshop [Gut09].

xor [FJ13].

yes [GPV24].

Zémor [TNS20]. zero [BG16, DT17, SW07b]. zero-knowledge [SW07b].

References

AlMashrafi:2013:IMI

[ABD^{+}13] Mufeed Al Mashrafi, Harry Bartlett, Ed Dawson, Leonie Simpson, and Kenneth Koon-Ho Wong. Indirect message injection for MAC generation. Journal of Mathematical Cryptology, 7(3):253–277, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Alomair:2010:PPS

[ACP10] Basel Alomair, Andrew Clark, and Radha Poovendran. The power of primes: security of authentication based on a universal hash-function family. Journal of Mathematical Cryptology, 4(2):121–148, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Asghar:2014:SCG

[ADPS14] Hassan Jameel Asghar, Yvo Desmedt, Josef Pieprzyk, and Ron Steinfeld. A subexponential construction of graph coloring for multiparty computation. Journal of Mathematical Cryptology, 8(4):363–403, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Antal:2017:MCP

[AGH17] Eugen Antal, Otokar Grošek, and Peter Horak. On a mnemonic construction of permutations. Journal of Mathematical Cryptology, 11(1):45–53, 2017. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Azizi:2019:CIC

[AJZT19] Abdelmalek Azizi, Idriss Jerrari, Abdelkader Zekhnini, and Mohammed Talbi. Capitulation of the 2-ideal classes of type (2, 2, 2) of some quartic cyclic number fields. Journal of Mathematical Cryptology, 13 (1):27–??, March 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2019.13.issue-
1/jmc-2017-0037/jmc-2017-
0037.xml.

Asghar:2017:WIP

[AK17] Hassan Jameel Asghar and Mohamed Ali Kaafar. When are identification protocols with sparse challenges safe? The case of the Coskun and Herley attack. Journal of Mathematical Cryptology, 11(3):177–194, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Aravamuthan:2007:ATO

[AL07] Sarang Aravamuthan and Sachin Lodha. The average transmission overhead for broadcast encryption. Journal of Mathematical Cryptology, 1(4):373–384, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Attrapadung:2011:FEP

[AL11] Nuttapong Attrapadung and Benoît Libert. Functional encryption for public-attribute inner products: achieving constant-size ciphertexts with adaptive security or support for negation. Journal of Mathematical Cryptology, 5(2):115–158, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Alawatugoda:2017:LRK

[Ala17] Janaka Alawatugoda. On the leakage-resilient key exchange. Journal of Mathematical Cryptology, 11(4):215–269, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Aly:2007:ELC

[AMW07] Hassan Aly, Wilfried Meidl, and Arne Winterhof. On the kerror linear complexity of cyclotomic sequences. Journal of Mathematical Cryptology, 1(3):

283–296, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Araujo:2020:CHR

[AN20] Filipe Araujo and Samuel Neves. The circulant hash revisited. Journal of Mathematical Cryptology, 15(1):250–257, December 3, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2018-0054/html.

Anonymous:2013:M

[Ano13] Anonymous. Masthead. Journal of Mathematical Cryptology, 7(3):i, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

//www.degruyter.com/view/
j/jmc.2013.7.issue-3/jmc-
2013-masthead3/jmc-2013-masthead3.
xml.

Anonymous:2017:F

[Ano17] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 11(2):i–iv, June 2017. ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

//www.degruyter.com/view/
j/jmc.2017.11.issue-2/jmc-
2017-frontmatter2/jmc-2017-
frontmatter2.xml.

Anonymous:2018:Fa

[Ano18a] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 12(1):i–??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-1/jmc-2018-
frontmatter1/jmc-2018-frontmatter1.
xml.

Anonymous:2018:Fb

[Ano18b] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 12(2):i–iv, June 2018. ISSN 1862-2976 (print), 1862-2984 (electronic).

Anonymous:2018:Fc

[Ano18c] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 12(3):i–??, September 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-3/jmc-2018-
frontmatter3/jmc-2018-frontmatter3.
xml.

Anonymous:2018:F

[Ano18d] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 12(4):i–??, December 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-4/jmc-2018-
frontmatter4/jmc-2018-frontmatter4.
xml.

Anonymous:2019:Fa

[Ano19a] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 13(1):i–??, March 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.

                                        degruyter.com/view/j/jmc.
                                        2019.13.issue-1/jmc-2019-
                                        frontmatter1/jmc-2019-frontmatter1.
                                        xml.

Anonymous:2019:Fb

[Ano19b] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 13(2):i–??, June 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

                                        tronic). URL http://www.
                                        degruyter.com/view/j/jmc.
                                        2019.13.issue-2/jmc-2019-
                                        frontmatter2/jmc-2019-frontmatter2.
                                        xml.

Anonymous:2019:F

[Ano19c] Anonymous. Frontmatter. Journal of Mathematical Cryptology, 13(3–4):i–??, September 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http:/

                                        /www.degruyter.com/view/j/
                                        jmc.2019.13.issue-3-4/jmc-
                                        2019-frontmatter3-4/jmc-2019-
                                        frontmatter3-4.xml.

Anokhin:2020:PFF

[Ano20] Mikhail Anokhin. Pseudo-free families of computational universal algebras. Journal of Mathematical Cryptology, 15(1): 197–222, November 25, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2020-0014/html.

Alam:2015:CWD

[AÖY15] Bilal Alam, Ferruh Özbudak, and Oğuz Yayla. Classes of weak Dembowski–Ostrom polynomials for multivariate quadratic cryptosystems. Journal of Mathematical Cryptology, 9(1): 11–22, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Albrecht:2015:CHL

[APS15] Martin R. Albrecht, Rachel Player, and Sam Scott. On the concrete hardness of learning with errors. Journal of Mathematical Cryptology, 9(3): 169–203, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Accardi:2015:CSA

[AR15] Luigi Accardi and Massimo Regoli. On a class of strongly asymmetric PKA algorithms. Journal of Mathematical Cryptology, 9(3):151–159, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

AlMahri:2018:FFA

[ASB^{+}18] Hassan Qahur Al Mahri, Leonie Simpson, Harry Bartlett, Ed Dawson, and Kenneth Koon-Ho Wong. A fundamental flaw in the ++AE authenticated encryption mode. Journal of Mathematical Cryptology, 12(1): 37–??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2018.12.issue-
1/jmc-2016-0037/jmc-2016-
0037.xml.

Avanzi:2008:RLG

[ATW08] R. Avanzi, N. Thériault, and Z. Wang. Rethinking low genus hyperelliptic Jacobian arithmetic over binary fields: interplay of field arithmetic and explicit formulæ. Journal of Mathematical Cryptology, 2(3): 227–255, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Banegas:2018:DKE

[BBB^{+}18] Gustavo Banegas, Paulo S. L. M. Barreto, Brice Odilon Boidje, Pierre-Louis Cayrel, Gilbert Ndollane Dione, Kris Gaj, Cheikh Thiécoumba Gueye, Richard Haeussler, Jean Belo Klamti, Ousmane N’diaye, Duc Tri Nguyen, Edoardo Persichetti, and Jefferson E. Ricardini. DAGS: Key encapsulation using dyadic GS codes. Journal of Mathematical Cryptology, 12(4):221–??, December 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-4/jmc-2018-
0027/jmc-2018-0027.xml.

Bindel:2019:EHL

[BBGS19] Nina Bindel, Johannes Buchmann, Florian Göpfert, and Markus Schmidt. Estimation of the hardness of the learning with errors problem with a restricted number of samples. Journal of Mathematical Cryptology, 13(1): 47–??, March 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2019.13.issue-
1/jmc-2017-0040/jmc-2017-
0040.xml.

Bansal:2019:SSI

[BBP19] Tarun Kumar Bansal, Xavier Boyen, and Josef Pieprzyk. Signcryption schemes with insider security in an ideal permutation model. Journal of Mathematical Cryptology, 13(2): 117–150, June 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

http://www.degruyter.com/
view/j/jmc.2019.13.issue-
2/jmc-2018-0006/jmc-2018-
0006.xml.

Biasse:2020:TBC

[BBP^{+}20] Jean-François Biasse, Xavier Bonnetain, Benjamin Pring, André Schrottenloher, and William Youmans. A trade-off between classical and quantum circuit size for an attack against CSIDH. Journal of Mathematical Cryptology, 15(1):4–17, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0070/html.

Banegas:2020:DED

[BBPS20] Gustavo Banegas, Paulo S.

L. M. Barreto, Edoardo Persichetti, and Paolo Santini. Designing efficient dyadic operations for cryptographic applications. Journal of Mathematical Cryptology, 14(1):95– 109, June 19, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2015-0054/html.

Brown:2012:RWS

[BC12] Alyssa Brown and Thomas W. Cusick. Recursive weights for some Boolean functions. Journal of Mathematical Cryptology, 6(2):105–135, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        http://www.degruyter.com/
                                        view/j/jmc.2012.6.issue-2/
                                        jmc-2011-0020/jmc-2011-0020.
                                        xml.

Bootland:2020:EPC

[BCIV20] Carl Bootland, Wouter Castryck, Ilia Iliashenko, and Frederik Vercauteren. Efficiently processing complexvalued data in homomorphic encryption. Journal of Mathematical Cryptology, 14(1):55– 65, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2015-0051/html.

Blackburn:2009:CPK

[BCM09] Simon R. Blackburn, Carlos Cid, and Ciaran Mullan. Cryptanalysis of the MST_{3} public key cryptosystem. Journal of Mathematical Cryptology, 3(4): 321–338, 2009. CODEN ????

ISSN 1862-2976 (print), 1862-2984 (electronic).

Blackburn:2011:CTM

[BCM11] Simon R. Blackburn, Carlos Cid, and Ciaran Mullan. Cryptanalysis of three matrix-based key establishment protocols. Journal of Mathematical Cryptology, 5(2):159–168, 2011. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bootland:2020:FCP

[BCSV20] Carl Bootland, Wouter Castryck, Alan Szepieniec, and Frederik Vercauteren. A framework for cryptographic problems from linear algebra. Journal of Mathematical Cryptology, 14(1):202–217, July 21, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0032/html.

Broon:2021:ITH

[BDFM21] Fouazou Lontouo Perez Broon, Thinh Dang, Emmanuel Fouotsa, and Dustin Moody. Isogenies on twisted Hessian curves. Journal of Mathematical Cryptology, 15(1):345–358, March 16, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2020-0037/html.

Bos:2014:CBA

[BDJ14] Joppe W. Bos, Alina Dudeanu, and Dimitar Jetchev. Collision bounds for the additive Pollard rho algorithm for solving discrete logarithms. Journal of Mathematical Cryptology, 8 (1):71–92, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Baumeister:2012:ALS

[BdW12] Barbara Baumeister and Jan-Hendrik de Wiljes. Aperiodic logarithmic signatures. Journal of Mathematical Cryptology, 6 (1):21–37, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bourgeois:2009:AAN

[BF09] Gérald Bourgeois and Jean-Charles Faugère. Algebraic attack on NTRU using Witt vectors and Gröbner bases. Journal of Mathematical Cryptology, 3 (3):205–214, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Brier:2020:NNT

[BFJN20] Éric Brier, Houda Ferradi, Marc Joye, and David Naccache. New number-theoretic cryptographic primitives. Journal of Mathematical Cryptology, 14(1):224– 235, August 1, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2019-0035/html.

Bouillaguet:2012:FWK

[BFJT12] Charles Bouillaguet, Pierre-Alain Fouque, Antoine Joux, and Joana Treger. A family of weak keys in HFE and the corresponding practical key-recovery. Journal of Mathematical Cryptology, 5(3–4):247–275, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bettale:2009:HAS

[BFP09] Luk Bettale, Jean-Charles Faugère, and Ludovic Perret. Hybrid approach for solving multivariate systems over finite fields. Journal of Mathematical Cryptology, 3(3):177–197, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bianco:2016:CTZ

[BG16] Giulia Bianco and Elisa Gorla.

Compression for trace zero points on twisted Edwards curves. Journal of Mathematical Cryptology, 10(1):15–34, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Boura:2020:CCR

[BGGJ20] Christina Boura, Nicolas Gama, Mariya Georgieva, and Dimitar Jetchev. CHIMERA: Combining Ring-LWE-based fully homomorphic encryption schemes. Journal of Mathematical Cryptology, 14(1):316– 338, August 7, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0026/html.

Boneh:2020:MNI

[BGK^{+}20] Dan Boneh, Darren Glass, Daniel Krashen, Kristin Lauter, Shahed Sharif, Alice Silverberg, Mehdi Tibouchi, and Mark Zhandry. Multiparty noninteractive key exchange and more from isogenies on elliptic curves. Journal of Mathematical Cryptology, 14(1):5– 14, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2015-0047/html.

Basu:2008:CCE

[BGMP08] Riddhipratim Basu, Shirshendu Ganguly, Subhamoy Maitra, and Goutam Paul. A complete characterization of the evolution of RC4 pseudo random generation algorithm. Journal of Mathematical Cryptology, 2(3): 257–289, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bisson:2011:CER

[Bis11] Gaetan Bisson. Computing endomorphism rings of elliptic curves under the GRH. Journal of Mathematical Cryptology, 5 (2):101–113, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Blake:2009:ORS

[BK09] Ian F. Blake and Vladimir Kolesnikov. One-round secure comparison of integers. Journal of Mathematical Cryptology, 3(1):37–68, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Brown:2022:CM

[BKL22] Daniel R. L. Brown, Neal Koblitz, and Jason T. LeGrow. Cryptanalysis of “MAKE”. Journal of Mathematical Cryptology, 16(1):98–102, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/document/doi/
10.1515/jmc-2021-0016/html.

See [RS22].

Blackburn:2009:CCG

[Bla09] Simon R. Blackburn. Cryptanalysing the critical group: efficiently solving Biggs’s discrete logarithm problem. Journal of Mathematical Cryptology, 3(3): 199–203, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Blackburn:2010:DLP

[Bla10] Simon R. Blackburn. The discrete logarithm problem modulo one: cryptanalysing the Ariffin– Abu cryptosystem. Journal of Mathematical Cryptology, 4(2): 193–198, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Bhattacharjee:2021:OVF

[BLLN21] Arghya Bhattacharjee, Cuauhtemoc Mancillas López, Eik List, and Mridul Nandi. The Oribatida v1.3 family of lightweight authenticated encryption schemes.

Journal of Mathematical Cryptology, 15(1):305–344, January 29, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0018/html.

Bockle:2009:MRT

[Böc09] Gebhard Böckle. The Miller– Rabin test with randomized exponents. Journal of Mathematical Cryptology, 3(4):307–319, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Boxall:2012:HPF

[Box12] John Boxall. Heuristics on pairing-friendly elliptic curves. Journal of Mathematical Cryptology, 6(2):81–104, 2012. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2012.6.issue-2/jmc-2011-0004/
jmc-2011-0004.xml?format=
INT.

Biasse:2020:FRO

[BP20] Jean-François Biasse and Benjamin Pring. A framework for reducing the overhead of the quantum oracle for use with Grover’s algorithm with applications to cryptanalysis of SIKE. Journal of Mathematical Cryptology, 15(1):143–156, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0080/html.

Basso:2022:SGA

[BP22] Andrea Basso and Fabien Pazuki. On the supersingular GPST attack. Journal of Mathematical Cryptology, 16(1): 14–19, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2021-0020/html.

Brown:2023:PKA

[Bro23] Daniel R. L. Brown. Plactic key agreement (insecure?). Journal of Mathematical Cryptology, 17 (1):??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2022-0010/html.

Blake:2007:SDC

[BS07] Ian F. Blake and Igor E. Shparlinski. Statistical distribution and collisions of VSH. Journal of Mathematical Cryptology, 1 (4):329–349, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Biasse:2019:QAA

[BS19] Jean-François Biasse and Fang Song. On the quantum attacks against schemes relying on the hardness of finding a short generator of an ideal in Q(ζ2s). Journal of Mathematical Cryptology, 13(3–4):151–??, September 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http:/

/www.degruyter.com/view/j/
jmc.2019.13.issue-3-4/jmc-
2015-0046/jmc-2015-0046.xml.

Banin:2012:DLP

[BT12] Matan Banin and Boaz Tsaban. The discrete logarithm problem in Bergman’s nonrepresentable ring. Journal of Mathematical Cryptology, 6(2): 171–182, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

//www.degruyter.com/view/
j/jmc.2012.6.issue-2/jmc-
2012-0014/jmc-2012-0014.xml.

Brandstatter:2009:ELC

[BW09] Nina Brandstätter and Arne Winterhof. k-error linear complexity over Fp of subsequences of Sidelnikov sequences of period (p^{r} − 1)/3. Journal of Mathematical Cryptology, 3(3): 215–225, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Chakraborty:2020:NAP

[CAR20] Suvradip Chakraborty, Janaka Alawatugoda, and Chandrasekaran Pandu Rangan. New approach to practical leakage-resilient public-key cryptography. Journal of Mathematical Cryptology, 14(1):172– 201, July 11, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/

document/doi/10.1515/jmc-
2019-0014/html.

Choi:2007:CHP

[CBW07] Su-Jeong Choi, Simon R. Blackburn, and Peter R. Wild. Cryptanalysis of a homomorphic public-key cryptosystem over a finite group. Journal of Mathematical Cryptology, 1(4):351– 358, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cusick:2015:TRS

[CC15] Thomas W. Cusick and Younhwan Cheon. Theory of 3-rotation symmetric cubic Boolean functions. Journal of Mathematical Cryptology, 9(1): 45–62, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cheon:2020:ACV

[CCH^{+}20] Jung Hee Cheon, Wonhee Cho, Minki Hhan, Jiseung Kim, and Changmin Lee. Algorithms for CRT-variant of approximate greatest common divisor problem. Journal of Mathematical Cryptology, 14(1):397–413, October 20, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0031/html.

Calderini:2024:DEU

[CCI24] Marco Calderini, Roberto Civino, and Riccardo Invernizzi. Differential experiments using parallel alternative operations. Journal of Mathematical Cryptology, 18 (1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2023-0030/html.

Carlet:2014:LSO

[CDGM14] Claude Carlet, Jean-Luc Danger, Sylvain Guilley, and Houssem Maghrebi. Leakage squeezing: optimal implementation and security evaluation. Journal of Mathematical Cryptology, 8(3):249–295, 2014. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Chatterjee:2018:CPB

[CDK18] Sanjit Chatterjee, M. Prem Laxman Das, and R. Kabaleeshwaran. Converting pairingbased cryptosystems from composite to prime order setting — a comparative analysis. Journal of Mathematical Cryptology, 12(3):159–??, September 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-3/jmc-2017-
0042/jmc-2017-0042.xml.

Chakraborti:2018:ONL

[CDN18] Avik Chakraborti, Nilanjan Datta, and Mridul Nandi. On the optimality of non-linear computations for symmetric key primitives. Journal of Mathematical Cryptology, 12(4):241– ??, December 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2018.12.issue-
4/jmc-2017-0011/jmc-2017-
0011.xml.

Castryck:2020:HFS

[CDS20] Wouter Castryck, Thomas Decru, and Benjamin Smith. Hash functions from superspecial genus-2 curves using Richelot isogenies. Journal of Mathematical Cryptology, 14(1):268– 292, August 7, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0021/html.

Chen:2015:SSK

[CEM15] Jiageng Chen, Keita Emura, and Atsuko Miyaji. SKENO: Secret key encryption with noninteractive opening. Journal of Mathematical Cryptology, 9 (2):63–74, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Coron:2020:ICA

[CG20] Jean-Sébastien Coron and Agnese Gini. Improved cryptanalysis of the AJPS Mersenne based cryptosystem. Journal of Mathematical Cryptology, 14(1):218– 223, July 21, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0027/html.

Chen:2013:SCM

[CGK13] Yanling Chen, Danilo Gligoroski, and Svein J. Knapskog. On a special class of multivariate quadratic quasigroups (MQQs). Journal of Mathematical Cryptology, 7(2):111–141, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Childs:2014:QCD

[CI14] Andrew M. Childs and Gábor Ivanyos. Quantum computation of discrete logarithms in semigroups. Journal of Mathematical Cryptology, 8(4):405–416, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Childs:2014:CEC

[CJS14] Andrew Childs, David Jao, and Vladimir Soukharev. Constructing elliptic curve isogenies in quantum subexponential time. Journal of Mathematical Cryptology, 8(1):1–29, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Carlet:2016:CPM

[CJST16] Claude Carlet, David Joyner, Pantelimon Stănică, and Deng Tang. Cryptographic properties of monotone Boolean functions. Journal of Mathematical Cryptology, 10(1):1–14, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

                                        tronic). URL http://www.
                                        degruyter.com/view/j/jmc.
                                        2016.10.issue-1/jmc-2014-
                                        0030/jmc-2014-0030.xml.

Colo:2020:OSI

[CK20] Leonardo Colò and David Kohel. Orienting supersingular isogeny graphs. Journal of Mathematical Cryptology, 14(1): 414–437, October 23, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0034/html.

Colbourn:2009:RCP

[CL09] Charles J. Colbourn and Alan C. H. Ling. A recursive construction for perfect hash families. Journal of Mathematical Cryptology, 3(4):291–306, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cusick:2009:CBS

[CLS09] Thomas W. Cusick, Yuan Li, and Pantelimon Stănică. On a conjecture for balanced symmetric Boolean functions. Journal of Mathematical Cryptology, 3 (4):273–290, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cusick:2016:AEM

[CLS16] Thomas W. Cusick, K. V.

Lakshmy, and M. Sethumadhavan. Affine equivalence of monomial rotation symmetric Boolean functions: a Pólya’s theorem approach. Journal of Mathematical Cryptology, 10(3– 4):145–156, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cheon:2020:EPS

[CLS20] Jung Hee Cheon, Kristin Lauter, and Yongsoo Song. Editor’s preface for the Second Annual MathCrypt Proceedings volume. Journal of Mathematical Cryptology, 15(1):1–3, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0170/html.

Calderini:2024:SER

[CLSV24] Marco Calderini, Riccardo Longo, Massimiliano Sala, and Irene Villa. Searchable encryption with randomized ciphertext and randomized keyword search. Journal of Mathematical Cryptology, 18(1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2023-0029/html.

Cheon:2020:PFA

[CLY20] Jung Hee Cheon, Kristin Lauter, and Donggeon Yhee. Preface to the First Annual MathCrypt Proceedings volume. Journal of Mathematical Cryptology, 14(1):1– 4, June 19, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0060/html.

Chakraborty:2022:MMA

[CN22] Bishwajit Chakraborty and Mridul Nandi. The mF mode of authenticated encryption with associated data. Journal of Mathematical Cryptology, 16(1): 73–97, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0054/html.

Cianfriglia:2024:MLC

[COP24] Marco Cianfriglia, Elia Onofri, and Marco Pedicini. mR_{LWE}- CP-ABE: A revocable CP-ABE for post-quantum cryptography. Journal of Mathematical Cryptology, 18(1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2023-0026/html.

Ciet:2011:CEC

[CQS11] Mathieu Ciet, Jean-Jacques Quisquater, and Francesco Sica. Compact elliptic curve representations. Journal of Mathematical Cryptology, 5(1):89–100, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Cascudo:2021:NSM

[CS21] Ignacio Cascudo and Reto Schnyder. A note on secure multiparty computation via higher residue symbols. Journal of Mathematical Cryptology, 15(1):

284–297, January 29, 2021. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2020-0013/html.

Chen:2024:DHS

[CS24] Imin Chen and David Sun. The dihedral hidden subgroup problem. Journal of Mathematical Cryptology, 18(1): ??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2022-0029/html.

Csirmaz:2020:SSD

[Csi20] Laszlo Csirmaz. Secret sharing and duality. Journal of Mathematical Cryptology, 15(1):157– 173, November 25, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0045/html.

Craven:2022:EGT

[CW22] Matthew J. Craven and John R. Woodward. Evolution of grouptheoretic cryptology attacks using hyper-heuristics. Journal of Mathematical Cryptology, 16(1): 49–63, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2021-0017/html.

Chee:2012:BCE

[CWZ12] Yeow Meng Chee, Huaxiong Wang, and Liang Feng Zhang. On the Bringer–Chabanne EPIR protocol for polynomial evaluation. Journal of Mathematical Cryptology, 5(3–4):277– 301, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Dibert:2014:ISS

[DC14] Alexander Dibert and László Csirmaz. Infinite secret sharing — examples. Journal of Mathematical Cryptology, 8(2): 141–168, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Duquesne:2014:ECP

[DEF14] Sylvain Duquesne, Nadia El Mrabet, and Emmanuel Fouotsa. Efficient computation of pairings on Jacobi quartic elliptic curves. Journal of Mathematical Cryptology, 8(4):331–362, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Dolev:2015:TEP

[DGG^{+}15] Shlomi Dolev, Juan Garay, Niv Gilboa, Vladimir Kolesnikov, and Yelena Yuditsky. Towards efficient private distributed computation on unbounded input streams. Journal of Mathematical Cryptology, 9(2):79–94, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

DeMicheli:2020:CON

[DHS20] Gabrielle De Micheli, Nadia Heninger, and Barak Shani. Characterizing overstretched NTRU attacks. Journal of Mathematical Cryptology, 14(1): 110–119, June 14, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2015-0055/html.

DeFeo:2014:TQR

[DJP14] Luca De Feo, David Jao, and Jérôme Plût. Towards quantumresistant cryptosystems from supersingular elliptic curve isogenies. Journal of Mathematical Cryptology, 8(3):209–247, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

DiCrescenzo:2020:DPG

[DKKS20] Giovanni Di Crescenzo, Matluba Khodjaeva, Delaram Kahrobaei, and Vladimir Shpilrain. Delegating a product of group exponentiations with application to signature schemes (submission to special NutMiC 2019 issue of JMC). Journal of Mathematical Cryptology, 14(1):438– 459, October 30, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0036/html.

Delaplace:2020:CWB

[DM20] Claire Delaplace and Alexander May. Can we beat the square root bound for ECDLP over F_{p^{2}} via representation? Journal of Mathematical Cryptology, 14(1): 293–306, August 18, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0025/html.

Demirkiran:2008:CHC

[DN08] Cevahir Demirkiran and Enric Nart. Counting hyperelliptic curves that admit a Koblitz model. Journal of Mathematical Cryptology, 2(2):163–179, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Dowden:2015:SMT

[Dow15] Chris Dowden. Secure message transmission in the presence of a fully generalised adversary. Journal of Mathematical Cryptology, 9(4):205–214, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Daemen:2007:PDC

[DR07] Joan Daemen and Vincent Rijmen. Probability distributions of correlation and differentials in block ciphers. Journal of Mathematical Cryptology, 1(3): 221–242, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Dehkordi:2017:CCG

[DS17] Massoud Hadian Dehkordi and Ali Safi. The complexity of the connected graph access structure on seven participants. Journal of Mathematical Cryptology, 11(1):25–35, 2017. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Dey:2018:GRB

[DS18] Sabyasachi Dey and Santanu Sarkar. Generalization of Roos bias in RC4 and some results on key–keystream relations. Journal of Mathematical Cryptology, 12(1):43–??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-1/jmc-2016-
0061/jmc-2016-0061.xml.

Doroz:2020:FNE

[DS20] Yarkin Doröz and Berk Sunar. Flattening NTRU for evaluation key free homomorphic encryption. Journal of Mathematical Cryptology, 14(1):66– 83, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2015-0052/html.

Dugardin:2021:SMD

[DSG21] Margaux Dugardin, Werner Schindler, and Sylvain Guilley. Stochastic methods defeat regular RSA exponentiation algorithms with combined blinding methods. Journal of Mathematical Cryptology, 15(1):408– 433, April 20, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0010/html.

Dachman-Soled:2020:SRL

[DSGKS20a] Dana Dachman-Soled, Huijing Gong, Mukul Kulkarni, and Aria Shahverdi. (In)Security of Ring-LWE under partial key exposure. Journal of Mathematical Cryptology, 15(1):72–86, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0075/html.

Dachman-Soled:2020:TRA

[DSGKS20b] Dana Dachman-Soled, Huijing Gong, Mukul Kulkarni, and Aria Shahverdi. Towards a ring analogue of the leftover hash lemma. Journal of Mathematical Cryptology, 15(1):87–110, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0076/html.

DiScala:2020:CNV

[DSS20] Antonio J. Di Scala, Carlo Sanna, and Edoardo Signorini. On the condition number of the Vandermonde matrix of the n th cyclotomic polynomial. Journal of Mathematical Cryptology, 15(1):174–178, November 25, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2020-0009/html.

Dehkordi:2017:MDZ

[DT17] Massoud Hadian Dehkordi and Roghayeh Taghizadeh. Multiple differential-zero correlation linear cryptanalysis of reducedround CAST-256. Journal of Mathematical Cryptology, 11(2): 55–62, June 2017. ISSN 1862-2976 (print), 1862-2984 (elec-

                                        tronic). URL https://www.
                                        degruyter.com/view/j/jmc.
                                        2017.11.issue-2/jmc-2016-
                                        0054/jmc-2016-0054.xml.

Duquesne:2011:RAA

[Duq11] Sylvain Duquesne. RNS arithmetic in Fpk and application to fast pairing computation. Journal of Mathematical Cryptology, 5(1):51–88, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Erdelyi:2024:LCR

[EHKN24] Márton Erdélyi, Pál Hegedüs, Sándor Z. Kiss, and Gábor P. Nagy. On linear codes with random multiplier vectors and the maximum trace dimension property. Journal of Mathematical Cryptology, 18(1): ??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2023-0022/html.

Ekeraa:2021:QAC

[Eke21] Martin Ekerå. Quantum algorithms for computing general discrete logarithms and orders with tradeoffs. Journal of Mathematical Cryptology, 15(1):359– 407, April 22, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0006/html.

Engelbert:2007:SMT

[EOS07] D. Engelbert, R. Overbeck, and A. Schmidt. A summary of McEliece-type cryptosystems and their security. Journal of Mathematical Cryptology, 1(2): 151–199, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Eskeland:2024:CPP

[Esk24] Sigurd Eskeland. Cryptanalysis of a privacy-preserving authentication scheme based on private set intersection. Journal of Mathematical Cryptology, 18 (1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2023-0032/html.

Ferradi:2020:RSP

[FGG^{+}20] Houda Ferradi, Rémi Géraud, Sylvain Guilley, David Naccache, and Mehdi Tibouchi. Recovering secrets from prefixdependent leakage. Journal of Mathematical Cryptology, 14(1):

15–24, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2015-0048/html.

Fuji-Hara:2008:TCM

[FHLMW08] Ryoh Fuji-Hara, Xiyang Li, Ying Miao, and Dianhua Wu. A TWOOA construction for multireceiver multi-message authentication codes. Journal of Mathematical Cryptology, 2(1):9–28, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Field:2013:UCT

[FJ13] Rebecca E. Field and Brant C. Jones. Using carry-truncated addition to analyze add-rotate-xor hash algorithms. Journal of Mathematical Cryptology, 7 (2):97–110, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Fotiadis:2018:GPF

[FK18] Georgios Fotiadis and Elisavet Konstantinou. Generating pairing-friendly elliptic curve parameters using sparse families. Journal of Mathematical Cryptology, 12(2):83–99, June 2018. ISSN 1862-2976 (print), 1862-2984 (electronic).

Fancsali:2008:SAF

[FL08] Sz. L. Fancsali and P. Ligeti. Some applications of finite geometry for secure network coding. Journal of Mathematical Cryptology, 2(3):209–225, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Flori:2013:ECF

[FM13] Jean-Pierre Flori and Sihem Mesnager. An efficient characterization of a family of hyperbent functions with multiple trace terms. Journal of Mathematical Cryptology, 7(1):43–68, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Fischer:2009:SRF

[FMS09] Simon Fischer, Willi Meier, and Dirk Stegemann. Some remarks on FCSRs and implications for stream ciphers. Journal of Mathematical Cryptology, 3 (3):227–236, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Frougny:2008:MWE

[FS08] Christiane Frougny and Wolfgang Steiner. Minimal weight expansions in Pisot bases. Journal of Mathematical Cryptology, 2(4):365–392, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Friedlander:2009:DSS

[FS09] John B. Friedlander and Igor E.

Shparlinski. On the density of some special primes. Journal of Mathematical Cryptology, 3(3): 265–271, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Fischlin:2011:SBS

[FS11] Marc Fischlin and Dominique Schröder. Security of blind signatures under aborts and applications to adaptive oblivious transfer. Journal of Mathematical Cryptology, 5(2):169–203, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Farashahi:2009:HEC

[FSV09] Reza R. Farashahi, Igor E. Shparlinski, and José Felipe Voloch. On hashing into elliptic curves. Journal of Mathematical Cryptology, 3(4):353–360, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Gallant:2012:FDL

[Gal12] Robert P. Gallant. Finding discrete logarithms with a set orbit distinguisher. Journal of Mathematical Cryptology, 6(1):1–20, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Gaudry:2007:FGA

[Gau07] P. Gaudry. Fast genus 2 arithmetic based on theta functions. Journal of Mathematical Cryptology, 1(3):243–265, 2007. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Gharahi:2013:PSS

[GD13] Motahhareh Gharahi and Massoud Hadian Dehkordi. Perfect secret sharing schemes for graph access structures on six participants. Journal of Mathematical Cryptology, 7(2):143–146, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Goel:2020:SSA

[GGD20] Neha Goel, Indivar Gupta, and B. K. Dass. Survey on SAP and its application in publickey cryptography. Journal of Mathematical Cryptology, 14(1): 144–152, July 3, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2016-0004/html.

Garber:2015:LBA

[GKL15] David Garber, Delaram Kahrobaei, and Ha T. Lam. Length-based attacks in polycyclic groups. Journal of Mathematical Cryptology, 9(1):33–43, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Galbraith:2007:SPC

[GÓS07] Steven D. Galbraith, Colm Ó hÉigeartaigh, and Caroline Sheedy. Simplified pairing computation and security implications. Journal of Mathematical Cryptology, 1(3):267–281, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Gaal:2009:SNE

[GP09] István Gaál and Michael E.

Pohst. On solving norm equations in global function fields.

Journal of Mathematical Cryptology, 3(3):237–248, 2009. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Grosek:2013:CS

[GP13] Otokar Grošek and Štefan Porubský. Coprime solutions to ax ≡ b (mod n). Journal of Mathematical Cryptology, 7(3): 217–224, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Gupta:2017:ADT

[GPR17] Kishan Chand Gupta, Sumit Kumar Pandey, and Indranil Ghosh Ray. Applications of design theory for the constructions of MDS matrices for lightweight cryptography. Journal of Mathematical Cryptology, 11(2):85– 116, June 2017. ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2017.11.issue-2/jmc-2016-
0013/jmc-2016-0013.xml.

Galbraith:2009:DMS

[GPRS09] Steven D. Galbraith, Jordi Pujolàs, Christophe Ritzenthaler, and Benjamin Smith. Distortion maps for supersingular genus two curves. Journal of Mathematical Cryptology, 3(1):1–18, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Ghantous:2024:ESB

[GPV24] Wissam Ghantous, Federico Pintore, and Mattia Veroni. Efficiency of SIDH-based signatures (yes, SIDH). Journal of Mathematical Cryptology, 18(1): ??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2023-0023/html.

Grzeskowiak:2020:VLS

[Grz20] Maciej Grześkowiak. A variant of the large sieve inequality with explicit constants. Journal of Mathematical Cryptology, 14(1):307–315, August 7, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0022/html.

Gutierrez:2009:FSW

[Gut09] Jaime Gutierrez. Foreword: Second Workshop on Mathematical Cryptology. Journal of Mathematical Cryptology, 3(3): 175–176, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). Held in Santander, October 23–25, 2008.

Galbraith:2013:SPH

[GZ13] Steven D. Galbraith and Chang-An Zhao. Self-pairings on hyperelliptic curves. Journal of Mathematical Cryptology, 7(1):31–42, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). See erratum [GZ14].

Galbraith:2014:ESP

[GZ14] Steven D. Galbraith and Chang-An Zhao. Erratum: Selfpairings on hyperelliptic curves [J. Math. Cryptol. 7 (2013), 31–42] [MR3101014]. Journal of Mathematical Cryptology, 8 (1):93, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). See [GZ13].

Harayama:2007:WSB

[HF07] Tomohiro Harayama and Donald K. Friesen. Weil sum for birthday attack in multivariate quadratic cryptosystem. Journal of Mathematical Cryptology, 1(1):79–104, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hinek:2008:SMP

[Hin08] M. Jason Hinek. On the security of multi-prime RSA. Journal of Mathematical Cryptology, 2(2): 117–147, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hitt:2009:FGC

[Hit09] Laura Hitt. Families of genus 2 curves with small embedding degree. Journal of Mathematical Cryptology, 3(1):19–36, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Huang:2020:QSP

[HKP^{+}20] Ming-Deh Huang, Michiel Kosters, Christophe Petit, Sze Ling Yeo, and Yang Yun. Quasi-subfield polynomials and the elliptic curve discrete logarithm problem. Journal of Mathematical Cryptology, 14(1):

25–38, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2015-0049/html.

Hinek:2009:ALS

[HL09] M. Jason Hinek and Charles C. Y. Lam. Another look at some fast modular arithmetic methods. Journal of Mathematical Cryptology, 3(2):165–174, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hinek:2010:CMA

[HL10] M. Jason Hinek and Charles C. Y. Lam. Common modulus attacks on small private exponent RSA and some fast variants (in practice). Journal of Mathematical Cryptology, 4(1):57–93, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Harsanyi:2019:EIR

[HL19] Károly Harsányi and Péter Ligeti. Exact information ratios for secret sharing on small graphs with girth at least 5. Journal of Mathematical Cryptology, 13(2):107–??, June 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL http://www.
degruyter.com/view/j/jmc.
2019.13.issue-2/jmc-2018-
0024/jmc-2018-0024.xml.

Heuberger:2007:MWC

[HM07] Clemens Heuberger and James A. Muir. Minimal weight and colexicographically minimal integer representations. Journal of Mathematical Cryptology, 1(4): 297–328, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hartung:2008:ISB

[HS08] Rupert J. Hartung and Claus-Peter Schnorr. Identification and signatures based on NPhard problems of indefinite quadratic forms. Journal of Mathematical Cryptology, 2(4): 327–341, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hameed:2015:CIW

[HS15a] Ali Hameed and Arkadii Slinko. A characterisation of ideal weighted secret sharing schemes. Journal of Mathematical Cryptology, 9(4):227–244, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Henry:2015:LAR

[HS15b] Kevin J. Henry and Douglas R. Stinson. Linear approaches to resilient aggregation in sensor networks. Journal of Mathematical Cryptology, 9(4):245–272, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hakuta:2010:EAS

[HST10] Keisuke Hakuta, Hisayoshi Sato, and Tsuyoshi Takagi. Efficient arithmetic on subfield elliptic curves over small finite fields of odd characteristic. Journal of Mathematical Cryptology, 4(3): 199–238, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hoffstein:2020:SSF

[HSWZ20] Jeffrey Hoffstein, Joseph H. Silverman, William Whyte, and Zhenfei Zhang. A signature scheme from the finite field isomorphism problem. Journal of Mathematical Cryptology, 14(1): 39–54, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2015-0050/html.

Haridas:2014:SAM

[HVV14] Deepthi Haridas, Sarma Venkataraman, and Geeta Varadan. Security analysis of modified Rivest scheme. Journal of Mathematical Cryptology, 8(3):297–303, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Hisil:2011:EAG

[HWCD11] Huseyin Hisil, Kenneth Koon-Ho Wong, Gary Carter, and Ed Dawson. An exploration of affine group laws for elliptic curves. Journal of Mathematical Cryptology, 5(1):1–50, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Isobe:2023:CEP

[IK23] Shuji Isobe and Eisuke Koizumi. A construction of encryption protocols over some semidirect products. Journal of Mathematical Cryptology, 17(1): ??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2022-0018/html.

Jirakitpuwapat:2018:NMC

[JCK^{+}18] Wachirapong Jirakitpuwapat, Parin Chaipunya, Poom Kumam, Sompong Dhompongsa, and Phatiphat Thounthong. New methods of construction of Cartesian authentication codes from geometries over finite commutative rings. Journal of Mathematical Cryptology, 12 (3):119–??, September 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

                                        tronic). URL https://www.
                                        degruyter.com/view/j/jmc.
                                        2018.12.issue-3/jmc-2017-
                                        0057/jmc-2017-0057.xml.

Jiang:2014:PAP

[Jia14] Shaoquan Jiang. Persistent asymmetric password-based key exchange. Journal of Mathematical Cryptology, 8(1):31–70, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Jao:2020:STP

[JLLRL20] David Jao, Jason LeGrow, Christopher Leonardi, and Luis Ruiz-Lopez. A subexponentialtime, polynomial quantum space algorithm for inverting the CM group action. Journal of Mathematical Cryptology, 14(1): 129–138, June 14, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2015-0057/html.

Joye:2020:EPR

[JLNN20] Marc Joye, Oleksandra Lapiha, Ky Nguyen, and David Naccache. The eleventh power residue symbol. Journal of Mathematical Cryptology, 15(1): 111–122, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2020-0077/html.

Jha:2016:RSG

[JN16] Ashwin Jha and Mridul Nandi.

Revisiting structure graphs: applications to CBC–MAC and EMAC. Journal of Mathematical Cryptology, 10(3–4):157–180, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Joye:2020:PEA

[Joy20] Marc Joye. Protecting ECC against fault attacks: The ring extension method revisited. Journal of Mathematical Cryptology, 14(1):254–267, August 1, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2019-0030/html.

Joux:2020:PNT

[JP20] Antoine Joux and Jacek Pomykała. Preface for the Number-Theoretic Methods in Cryptology conferences. Journal of Mathematical Cryptology, 14(1):393–396, October 13, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2019-0111/html.

Jhanwar:2013:USI

[JSN13] Mahabir P. Jhanwar and Reihaneh Safavi-Naini. Unconditionallysecure ideal robust secret sharing schemes for threshold and multilevel access structure. Journal of Mathematical Cryptology, 7(4):279–296, 2013. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Justus:2014:DQR

[Jus14] Benjamin Justus. The distribution of quadratic residues and non-residues in the Goldwasser– Micali type of cryptosystem. Journal of Mathematical Cryptology, 8(2):115–140, 2014. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Justus:2015:DQR

[Jus15] Benjamin Justus. The distribution of quadratic residues and non-residues in the Goldwasser– Micali type of cryptosystem. II.

Journal of Mathematical Cryptology, 9(2):115–137, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Wei:2012:FMP

[jWW12] Tzer jen Wei and Lih-Chung Wang. A fast mental poker protocol. Journal of Mathematical Cryptology, 6(1):39–68, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Karabina:2010:FCC

[Kar10] Koray Karabina. Factor-4 and 6 compression of cyclotomic subgroups of F24m^{∗} and F36m^{∗}. Journal of Mathematical Cryptology, 4(1):1–42, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Karmakar:2020:EPF

[Kar20a] Sudhangshu B. Karmakar. An elementary proof of Fermat’s Last Theorem for all even exponents. Journal of Mathematical Cryptology, 14(1):139– 142, July 3, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2016-0018/html. See retraction

[Kar20b].

Karmakar:2020:REP

[Kar20b] Sudhangshu B. Karmakar. Retraction of: An elementary proof of Fermat’s Last Theorem for all even exponents. Journal of Mathematical Cryptology, 14 (1):143, July 3, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2017-2000/html. See [Kar20a].

Klamti:2023:CBH

[KH23] Jean Belo Klamti and M. Anwarul Hasan. A code-based hybrid signcryption scheme. Journal of Mathematical Cryptology, 17(1):??, January 2023. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2022-0002/html.

Kortelainen:2010:MAG

[KHK10] Juha Kortelainen, Kimmo Halunen, and Tuomas Kortelainen. Multicollision attacks and generalized iterated hash functions. Journal of Mathematical Cryptology, 4(3):239–270, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Kawa:2023:ASD

[KK23] Renata Kawa and Mieczysław Kula. Access structures determined by uniform polymatroids. Journal of Mathematical Cryptology, 17(1):??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2022-0017/html.

Koblitz:2008:ALN

[KM08] Neal Koblitz and Alfred Menezes.

Another look at non-standard discrete log and Diffie–Hellman problems. Journal of Mathematical Cryptology, 2(4):311–326, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Koblitz:2013:ALH

[KM13] Neal Koblitz and Alfred Menezes.

Another look at HMAC. Journal of Mathematical Cryptology, 7(3):225–251, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Kaji:2019:PEP

[KMNN19] Shizuo Kaji, Toshiaki Maeno, Koji Nuida, and Yasuhide Numata. Polynomial expressions of p-ary auction functions. Journal of Mathematical Cryptology, 13(2):69–??, June 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL http://www.
degruyter.com/view/j/jmc.
2019.13.issue-2/jmc-2018-
0016/jmc-2018-0016.xml.

Karabina:2010:AEW

[KMPS10] Koray Karabina, Alfred Menezes, Carl Pomerance, and Igor E. Shparlinski. On the asymptotic effectiveness of Weil descent attacks. Journal of Mathematical Cryptology, 4(2):175–191, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Kotov:2020:AKP

[KMU20] Matvei Kotov, Anton Menshov, and Alexander Ushakov. Attack on kayawood protocol: uncloaking private keys. Journal of Mathematical Cryptology, 15 (1):237–249, December 1, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2019-0015/html.

Kurosawa:2013:NLR

[KNP13] Kaoru Kurosawa, Ryo Nojima, and Le Trieu Phong. New leakage-resilient CCA-secure public key encryption. Journal of Mathematical Cryptology, 7 (4):297–312, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Koblitz:2007:ALA

[Kob07] Neal Koblitz. Another look at automated theorem-proving. Journal of Mathematical Cryptology, 1(4):385–403, 2007. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Koblitz:2012:ALA

[Kob12] Neal Koblitz. Another look at automated theorem-proving II. Journal of Mathematical Cryptology, 5(3–4):205–224, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Kim:2020:EAC

[KT20] Taechan Kim and Mehdi Tibouchi. Equidistribution among cosets of elliptic curve points in intervals. Journal of Mathematical Cryptology, 14(1):339– 345, August 7, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0020/html.

Kotov:2015:ACP

[KU15] Matvei Kotov and Alexander Ushakov. Analysis of a certain polycyclic-group-based cryptosystem. Journal of Mathematical Cryptology, 9(3):161– 167, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Kotov:2018:AKE

[KU18] Matvei Kotov and Alexander Ushakov. Analysis of a key exchange protocol based on tropical matrix algebra. Journal of Mathematical Cryptology, 12(3):137–??, September 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-3/jmc-2016-
0064/jmc-2016-0064.xml.

Kushwaha:2018:ILB

[Kus18] Prabhat Kushwaha. Improved lower bound for Diffie–Hellman problem using multiplicative group of a finite field as auxiliary group. Journal of Mathematical Cryptology, 12(2):101–118, June 2018. ISSN 1862-2976 (print), 1862-2984 (electronic).

Kousidis:2019:FFD

[KW19] Stavros Kousidis and Andreas Wiemers. On the first fall degree of summation polynomials. Journal of Mathematical Cryptology, 13(3–4):229–??, September 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http:/

                                        /www.degruyter.com/view/j/
                                        jmc.2019.13.issue-3-4/jmc-
                                        2017-0022/jmc-2017-0022.xml.

Laarhoven:2020:AVC

[Laa20] Thijs Laarhoven. Approximate Voronoi cells for lattices, revisited. Journal of Mathematical Cryptology, 15(1):60–71, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2020-0074/html.

Li:2007:SAC

[LC07] Yuan Li and T. W. Cusick. Strict avalanche criterion over finite fields. Journal of Mathematical Cryptology, 1(1):65–78, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Lavauzelle:2019:GCP

[LdV19] Julien Lavauzelle and Françoise Levy dit Vehel. Generic constructions of PoRs from codes and instantiations. Journal of Mathematical Cryptology, 13(2):81–??, June 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http://www.

degruyter.com/view/j/jmc.
2019.13.issue-2/jmc-2018-
0018/jmc-2018-0018.xml.

Laine:2015:TMT

[LL15] Kim Laine and Kristin Lauter.

Time-memory trade-offs for index calculus in genus 3. Journal of Mathematical Cryptology, 9(2):95–114, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Luykx:2015:TPB

[LMPW15] Atul Luykx, Bart Mennink, Bart Preneel, and Laura Winnen. Two-permutation-based hashing with binary mixing. Journal of Mathematical Cryptology, 9(3):139–150, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Lamberger:2009:NSC

[LNR09] Mario Lamberger, Tomislav Nad, and Vincent Rijmen. Numerical solvers and cryptanalysis. Journal of Mathematical Cryptology, 3(3):249–263, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Lu:2017:CRV

[LPS17] Yao Lu, Liqiang Peng, and Santanu Sarkar. Cryptanalysis of an RSA variant with moduli N = p^{r}q^{l}. Journal of Mathematical Cryptology, 11(2):117–130, June 2017. ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2017.11.issue-

2/jmc-2016-0025/jmc-2016-
0025.xml.

Lesavourey:2020:SPI

[LPS20] Andrea Lesavourey, Thomas Plantard, and Willy Susilo. Short principal ideal problem in multicubic fields. Journal of Mathematical Cryptology, 14(1): 359–392, August 20, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0028/html.

Lange:2007:DSS

[LS07] Tanja Lange and Igor E. Shparlinski. Distribution of some sequences of points on elliptic curves. Journal of Mathematical Cryptology, 1(1):1–11, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Laing:2018:SRR

[LS18] Thalia M. Laing and Douglas R. Stinson. A survey and refinement of repairable threshold schemes. Journal of Mathematical Cryptology, 12(1):57– ??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2018.12.issue-
1/jmc-2017-0058/jmc-2017-
0058.xml.

Longrigg:2008:CSC

[LU08] Jonathan Longrigg and Alexander Ushakov. Cryptanalysis of the shifted conjugacy authentication protocol. Journal of Mathematical Cryptology, 2(2): 109–116, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Liu:2023:FRR

[LZ23] Zhenyu Liu and Zepeng Zhuo.

Further research results on confusion coefficient of Boolean functions. Journal of Mathematical Cryptology, 17(1): ??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2021-0039/html.

Mohammed:2017:RNF

[MA17] Ahmed Mohammed and Abdulrahman Alkhelaifi. RSA: A number of formulas to improve the search for p + q. Journal of Mathematical Cryptology, 11(4): 195–203, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Magliveras:2013:F

[Mag13] Spyros S. Magliveras. Foreword. Journal of Mathematical Cryptology, 7(3):181–182, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

//www.degruyter.com/view/
j/jmc.2013.7.issue-3/jmc-
2013-5001/jmc-2013-5001.xml.

Maze:2012:AKD

[Maz12] Gérard Maze. Analysis of a key distribution scheme in secure multicasting. Journal of Mathematical Cryptology, 6(1):69–80, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Menezes:2007:ALH

[Men07] Alfred Menezes. Another look at HMQV. Journal of Mathematical Cryptology, 1(1):47–64, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Marti-Farre:2010:SSS

[MFP10] Jaume Martí-Farré and Carles Padró. On secret sharing schemes, matroids and polymatroids. Journal of Mathematical Cryptology, 4(2):95–120, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Mie:2008:PTR

[Mie08] Thilo Mie. Polylogarithmic tworound argument systems. Journal of Mathematical Cryptology, 2(4):343–363, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Martin:2007:CGC

[MN07] Keith Martin and Siaw-Lynn Ng. The combinatorics of generalised cumulative arrays. Journal of Mathematical Cryptology, 1(1):13–32, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Moch:2023:PSA

[Moc23] Alexander Moch. Provable security against generic attacks on stream ciphers. Journal of Mathematical Cryptology, 17(1): ??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2022-0033/html.

Morales:2008:ADE

[Mor08] David J. Mireles Morales. An attack on disguised elliptic curves. Journal of Mathematical Cryptology, 2(1):1–8, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Murphy:2008:GVC

[MP08] S. Murphy and M. B. Paterson. A geometric view of cryptographic equation solving. Journal of Mathematical Cryptology, 2(1):63–107, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Murphy:2020:DPD

[MP20] Sean Murphy and Rachel Player. Discretisation and product distributions in Ring-LWE. Journal of Mathematical Cryptology, 15(1):45–59, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0073/html.

Moody:2016:ISF

[MPST16] Dustin Moody, Souradyuti Paul, and Daniel Smith-Tone. Indifferentiability security of the fast wide pipe hash: breaking the birthday barrier. Journal of Mathematical Cryptology, 10(2):

101–133, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

McGuire:2020:LST

[MR20] Gary McGuire and Oisín Robinson. Lattice sieving in three dimensions for discrete log in medium characteristic. Journal of Mathematical Cryptology, 15(1):223–236, November 25, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0008/html.

Mouha:2012:CIR

[MSP12] Nicky Mouha, Gautham Sekar, and Bart Preneel. Challenging the increased resistance of regular hash functions against birthday attacks. Journal of Mathematical Cryptology, 6(3– 4):229–248, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2012.6.issue-3-
4/jmc-2011-0010/jmc-2011-
0010.xml.

Myasnikov:2008:RSA

[MU08] Alexei G. Myasnikov and Alexander Ushakov. Random subgroups and analysis of the length-based and quotient attacks. Journal of Mathematical Cryptology, 2(1):29–61, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Mosina:2010:MSA

[MU10] Natalia Mosina and Alexander Ushakov. Mean-set attack: cryptanalysis of Sibert et al. authentication protocol. Journal of Mathematical Cryptology, 4 (2):149–174, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Myasnikov:2014:CMC

[MU14] Alex D. Myasnikov and Alexander Ushakov. Cryptanalysis of matrix conjugation schemes. Journal of Mathematical Cryptology, 8(2):95–114, 2014. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Mullan:2011:CVS

[Mul11] Ciaran Mullan. Cryptanalysing variants of Stickel’s key agreement scheme. Journal of Mathematical Cryptology, 4(4):365– 373, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Murphy:2012:ELH

[Mur12] Sean Murphy. The effectiveness of the linear hull effect. Journal of Mathematical Cryptology, 6 (2):137–147, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

//www.degruyter.com/view/
j/jmc.2012.6.issue-2/jmc-
2011-0025/jmc-2011-0025.xml.

Moody:2012:FEC

[MW12] Dustin Moody and Hongfeng Wu. Families of elliptic curves with rational 3-torsion. Journal of Mathematical Cryptology, 5(3–4):225–246, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Nandi:2009:ISA

[Nan09] Mridul Nandi. Improved security analysis for OMAC as a pseudorandom function. Journal of Mathematical Cryptology, 3(2):133–148, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Nandi:2008:ISA

[NM08] Mridul Nandi and Avradip Mandal. Improved security analysis of PMAC. Journal of Mathematical Cryptology, 2(2): 149–162, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Nandi:2016:SJS

[NP16] Mridul Nandi and Tapas Pandit. On the security of joint signature and encryption revisited. Journal of Mathematical Cryptology, 10(3–4):181–221, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Nandi:2019:PSP

[NP19] Mridul Nandi and Tapas Pandit. Predicate signatures from pair encodings via dual system proof technique. Journal of Mathematical Cryptology, 13(3–4):197–??, September 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http:/

/www.degruyter.com/view/j/
jmc.2019.13.issue-3-4/jmc-
2017-0007/jmc-2017-0007.xml.

Neven:2009:HFR

[NSW09] Gregory Neven, Nigel P. Smart, and Bogdan Warinschi. Hash function requirements for Schnorr signatures. Journal of Mathematical Cryptology, 3(1):69–87, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Nguyen:2008:SAS

[NV08] Phong Q. Nguyen and Thomas Vidick. Sieve algorithms for the shortest vector problem are practical. Journal of Mathematical Cryptology, 2(2):181–207, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Okano:2012:VCF

[Oka12] Keiji Okano. On the ρ-values of complete families of pairingfriendly elliptic curves. Journal of Mathematical Cryptology, 6 (3–4):249–268, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2012.6.issue-3-
4/jmc-2012-0011/jmc-2012-
0011.xml?format=INT.

Orumiehchiha:2013:SAL

[OPSB13] Mohammad Ali Orumiehchiha, Josef Pieprzyk, Ron Steinfeld, and Harry Bartlett. Security analysis of linearly filtered NLF-SRs. Journal of Mathematical Cryptology, 7(4):313–332, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Omar:2014:FHF

[OS14] Sami Omar and Houssem Sabri. Fast hash functions and convolution product. Journal of Mathematical Cryptology, 8(2): 169–187, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Partala:2018:AGD

[Par18] Juha Partala. Algebraic generalization of Diffie–Hellman key exchange. Journal of Mathematical Cryptology, 12(1):1– ??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        view/j/jmc.2018.12.issue-
                                        1/jmc-2017-0015/jmc-2017-
                                        0015.xml.

Persichetti:2012:CMK

[Per12] Edoardo Persichetti. Compact McEliece keys based on quasidyadic Srivastava codes. Journal of Mathematical Cryptology, 6 (2):149–169, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

                                        //www.degruyter.com/view/
                                        j/jmc.2012.6.issue-2/jmc-
                                        2011-0099/jmc-2011-0099.xml.

Pandey:2020:ICE

[PGS20] Atul Pandey, Indivar Gupta, and Dhiraj Kumar Singh. Improved cryptanalysis of a ElGamal cryptosystem based on matrices over group rings. Journal of Mathematical Cryptology, 15 (1):266–279, December 20, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0054/html.

Popov:2017:DTP

[Pop17] Serguei Popov. On a decentralized trustless pseudo-random number generation algorithm. Journal of Mathematical Cryptology, 11(1):37–43, 2017. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Poulakis:2016:NLA

[Pou16] Dimitrios Poulakis. New lattice attacks on DSA schemes. Journal of Mathematical Cryptology, 10(2):135–144, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Pinto:2018:BPF

[PP18] Eduardo Carvalho Pinto and Christophe Petit. Better pathfinding algorithms in LPS Ramanujan graphs. Journal of Mathematical Cryptology, 12(4): 191–??, December 2018. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2018.12.issue-4/jmc-2017-
0051/jmc-2017-0051.xml.

Pomykala:2020:IFC

[PR20] Jacek Pomykała and Maciej Radziejewski. Integer factoring and compositeness witnesses. Journal of Mathematical Cryptology, 14(1):346–358, August 20, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2019-0023/html.

Paterson:2008:TAS

[PS08] M. B. Paterson and D. R. Stinson. Two attacks on a sensor network key distribution scheme of Cheng and Agrawal. Journal of Mathematical Cryptology, 2 (4):393–403, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Paterson:2015:OCI

[PS15] Maura B. Paterson and Douglas R. Stinson. Optimal constructions for ID-based oneway-function key predistribution schemes realizing specified communication graphs. Journal of Mathematical Cryptology, 9 (4):215–225, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Paterson:2020:EAC

[PS20] Maura B. Paterson and Douglas R. Stinson. On the equivalence of authentication codes and robust (2, 2)-threshold schemes. Journal of Mathematical Cryptology, 15(1):179–196, November 25, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/

document/doi/10.1515/jmc-
2019-0048/html.

Paterson:2013:CTF

[PSU13] Maura B. Paterson, Douglas R.

Stinson, and Jalaj Upadhyay. A coding theory foundation for the analysis of general unconditionally secure proof-of-retrievability schemes for cloud storage. Journal of Mathematical Cryptology, 7(3):183–216, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Paterson:2018:MPP

[PSU18] Maura B. Paterson, Douglas R.

Stinson, and Jalaj Upadhyay. Multi-prover proof of retrievability. Journal of Mathematical Cryptology, 12(4):203– ??, December 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2018.12.issue-
4/jmc-2018-0012/jmc-2018-
0012.xml.

Reichl:2017:TLS

[Rei17] Dominik Reichl. Tame logarithmic signatures of abelian groups. Journal of Mathematical Cryptology, 11(4):205–214, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Reid:2021:UIE

[Rei21] Elizabeth M. Reid. Using inclusion/exclusion to find bent and balanced monomial rotation symmetric functions. Journal of Mathematical Cryptology, 15 (1):298–304, January 29, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2020-0021/html.

Rudy:2020:RTK

[RM20] Dylan Rudy and Chris Monico. Remarks on a tropical key exchange system. Journal of Mathematical Cryptology, 15(1): 280–283, December 20, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0061/html.

Rouquette:2024:ABA

[RMS24] Loïc Rouquette, Marine Minier, and Christine Solnon. Automatic boomerang attacks search on Rijndael. Journal of Mathematical Cryptology, 18 (1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2023-0027/html.

Rahman:2022:MMA

[RS22] Nael Rahman and Vladimir Shpilrain. MAKE: a matrix action key exchange. Journal of Mathematical Cryptology, 16(1): 64–72, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-

2020-0053/html. See successful

attack [BKL22].

Ruinskiy:2007:LBC

[RST07] Dima Ruinskiy, Adi Shamir, and Boaz Tsaban. Lengthbased cryptanalysis: the case of Thompson’s group. Journal of Mathematical Cryptology, 1(4): 359–372, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Romankov:2023:AQA

[RUS23] Vitaly Roman’kov, Alexander Ushakov, and Vladimir Shpilrain. Algebraic and quantum attacks on two digital signature schemes. Journal of Mathematical Cryptology, 17(1): ??, January 2023. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2022-0023/html.

Raddum:2018:MSB

[RZ18] Håvard Raddum and Pavol Zajac. MRHS solver based on linear algebra and exhaustive search. Journal of Mathematical Cryptology, 12(3):143– ??, September 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2018.12.issue-
3/jmc-2017-0005/jmc-2017-
0005.xml.

Schindler:2008:ASM

[Sch08] Werner Schindler. Advanced stochastic methods in side channel analysis on block ciphers in the presence of masking. Journal of Mathematical Cryptology, 2(3):291–310, 2008. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Scholl:2017:IEC

[Sch17] Travis Scholl. Isolated elliptic curves and the MOV attack. Journal of Mathematical Cryptology, 11(3):131–146, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Sha:2014:NIC

[Sha14] Min Sha. On the non-idealness of cyclotomic families of pairingfriendly elliptic curves. Journal of Mathematical Cryptology, 8 (4):417–440, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Shokrieh:2010:MPD

[Sho10] Farbod Shokrieh. The monodromy pairing and discrete logarithm on the Jacobian of finite graphs. Journal of Mathematical Cryptology, 4(1):43–56, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Sica:2020:FH

[Sic20] Francesco Sica. Factoring with hints. Journal of Mathematical Cryptology, 15(1):123–130, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2020-0078/html.

Silverman:2007:OPS

[Sil07] Robert D. Silverman. Optimal parameterization of SNFS. Journal of Mathematical Cryptology, 1(2):105–124, 2007. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Santini:2022:RFC

[SPB22] Paolo Santini, Edoardo Persichetti, and Marco Baldi. Reproducible families of codes and cryptographic applications. Journal of Mathematical Cryptology, 16(1):20– 48, January 2022. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0003/html.

Sipasseuth:2019:EGG

[SPS19] Arnaud Sipasseuth, Thomas Plantard, and Willy Susilo. Enhancing Goldreich, Goldwasser and Halevi’s scheme with intersecting lattices. Journal of Mathematical Cryptology, 13(3–4):169–??, September 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL http:/

/www.degruyter.com/view/j/
jmc.2019.13.issue-3-4/jmc-
2016-0066/jmc-2016-0066.xml.

Sepahi:2012:NSN

[SPSS12] Reza Sepahi, Josef Pieprzyk, Siamak F. Shahandashti, and Berry Schoenmakers. New security notions and relations for public-key encryption. Journal of Mathematical Cryptology, 6 (3–4):183–227, 2012. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Saxena:2009:CPB

[SS09] Amitabh Saxena and Ben Soh. A cryptographic primitive based on hidden-order groups. Journal of Mathematical Cryptology, 3(2):89–132, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Samajder:2016:ALN

[SS16a] Subhabrata Samajder and Palash Sarkar. Another look at normal approximations in cryptanalysis. Journal of Mathematical Cryptology, 10(2):69–99, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Swanson:2016:USS

[SS16b] Colleen M. Swanson and Douglas R. Stinson. Unconditionally secure signature schemes revisited. Journal of Mathematical Cryptology, 10(1):35–67, 2016. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Samajder:2017:RUB

[SS17] Subhabrata Samajder and Palash Sarkar. Rigorous upper bounds on data complexities of block cipher cryptanalysis. Journal of Mathematical Cryptology, 11(3):147–175, 2017. ISSN 1862-2976 (print), 1862-2984 (electronic).

Saraswat:2017:SAP

[SSA17] Vishal Saraswat, Rajeev Anand Sahu, and Amit K. Awasthi. A secure anonymous proxy signcryption scheme. Journal of Mathematical Cryptology, 11(2): 63–84, June 2017. ISSN 1862-2976 (print), 1862-2984 (elec-

tronic). URL https://www.
degruyter.com/view/j/jmc.
2017.11.issue-2/jmc-2015-
0014/jmc-2015-0014.xml.

Singh:2011:PDE

[SSS11] Rajesh P. Singh, A. Saikia, and B. K. Sarma. Poly-dragon: an efficient multivariate public key cryptosystem. Journal of Mathematical Cryptology, 4(4): 349–364, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Sala:2024:GSE

[ST24] Massimiliano Sala and Daniele Taufer. Group structure of elliptic curves over Z/NZ. Journal of Mathematical Cryptology, 18(1):??, January 2024. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2023-0025/html.

Stinson:2014:EDS

[SU14] Douglas R. Stinson and Jalaj Upadhyay. Is extracting data the same as possessing data? Journal of Mathematical Cryptology, 8(2):189–207, 2014. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Svaba:2010:PKC

[SvT10] Pavol Svaba and Tran van Trung. Public key cryptosystem MST_{3}: cryptanalysis and realization. Journal of Mathematical Cryptology, 4(3):271–315, 2010. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Staszewski:2013:SAL

[SvT13] Reiner Staszewski and Tran van Trung. Strongly aperiodic logarithmic signatures. Journal of Mathematical Cryptology, 7(2): 147–179, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL https:

                                        //www.degruyter.com/view/
                                        j/jmc.2013.7.issue-2/jmc-
                                        2013-5000/jmc-2013-5000.xml.

Stinson:2007:SRQ

[SW07a] D. R. Stinson and R. Wei. Some results on query processes and reconstruction functions for unconditionally secure 2-server 1round binary private information retrieval protocols. Journal of Mathematical Cryptology, 1(1):33–46, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Stinson:2007:EST

[SW07b] D. R. Stinson and J. Wu. An efficient and secure two-flow zeroknowledge identification protocol. Journal of Mathematical Cryptology, 1(3):201–220, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Sparr:2015:RFK

[SW15] Rüdiger Sparr and Ralph Wernsdorf. The round functions of KASUMI generate the alternating group. Journal of Mathematical Cryptology, 9(1):23–32, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Timpanella:2024:GDL

[Tim24] Marco Timpanella. On a generalization of the Deligne–Lusztig curve of Suzuki type and application to AG codes. Journal of Mathematical Cryptology, 18 (1):??, January 2024. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2023-0024/html.

Tischhauser:2011:NCA

[Tis11]  Elmar Tischhauser.  Nons-

mooth cryptanalysis, with an application to the stream cipher MICKEY. Journal of Mathematical Cryptology, 4(4): 317–348, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Tassa:2013:OEM

[TJBY13] Tamir Tassa, Ayman Jarrous, and Yonatan Ben-Ya’akov. Oblivious evaluation of multivariate polynomials. Journal of Mathematical Cryptology, 7 (1):1–29, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Takahashi:2020:AAS

[TKF^{+}20] Yasushi Takahashi, Momonari Kudo, Ryoya Fukasaku, Yasuhiko Ikematsu, Masaya Yasuda, and Kazuhiro Yokoyama. Algebraic approaches for solving isogeny problems of prime power degrees. Journal of Mathematical Cryptology, 15(1):31– 44, November 17, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2020-0072/html.

Tsaban:2015:CMS

[TL15] Boaz Tsaban and Noam Lifshitz. Cryptanalysis of the MORE symmetric key fully homomorphic encryption scheme. Journal of Mathematical Cryptology, 9(2):75–78, 2015. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Tomkins:2020:NZT

[TNS20] Hayley Tomkins, Monica Nevins, and Hadi Salmasian. New Zémor–Tillich type hash functions over GL_{2}(F_{p^{n}}). Journal of Mathematical Cryptology, 14(1): 236–253, August 1, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2019-0033/html.

Taraskin:2020:TIB

[TSJL20] Oleg Taraskin, Vladimir Soukharev, David Jao, and Jason T.

LeGrow. Towards isogeny-based password-authenticated key establishment. Journal of Mathematical Cryptology, 15(1):18– 30, November 17, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2020-0071/html.

Tibouchi:2020:OBA

[TW20] Mehdi Tibouchi and Alexandre Wallet. One bit is all it takes: A devastating timing attack on BLISS’s non-constant time sign flips. Journal of Mathematical Cryptology, 15(1):131–142, November 17, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2020-0079/html.

Urbanik:2020:NTS

[UJ20] David Urbanik and David Jao.

New techniques for SIDH-based NIKE. Journal of Mathematical Cryptology, 14(1):120– 128, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
document/doi/10.1515/jmc-
2015-0056/html.

vanTrung:2018:CSA

[vT18] Tran van Trung. Construction of strongly aperiodic logarithmic signatures. Journal of Mathematical Cryptology, 12(1): 23–??, March 2018. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        view/j/jmc.2018.12.issue-
                                        1/jmc-2017-0048/jmc-2017-
                                        0048.xml.

vonzurGathen:2009:SSP

[vzGS09] Joachim von zur Gathen and Igor E. Shparlinski. Subset sum pseudorandom numbers: fast generation and distribution. Journal of Mathematical Cryptology, 3(2):149–163, 2009. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

vonzurGathen:2013:GSP

[vzGS13] Joachim von zur Gathen and Igor E. Shparlinski. Generating safe primes. Journal of Mathematical Cryptology, 7(4): 333–365, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Walker:2007:PHF

[WC07] Robert A. Walker II and Charles J. Colbourn. Perfect Hash families: constructions and existence. Journal of Mathematical Cryptology, 1(2): 125–150, 2007. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Wang:2013:MCN

[WL13] Tianze Wang and Dongdai Lin. A method for counting the number of polynomial equivalence classes. Journal of Mathematical Cryptology, 7(1):69–95, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Wolf:2011:EKU

[WP11] Christopher Wolf and Bart Preneel. Equivalent keys in Multivariate Quadratic public key systems. Journal of Mathematical Cryptology, 4(4): 375–415, 2011. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Wu:2009:EIP

[WS09] J. Wu and D. R. Stinson. An efficient identification protocol secure against concurrent-reset attacks. Journal of Mathematical Cryptology, 3(4):339–352, 2009. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Wunderer:2019:DAH

[Wun19]  Thomas Wunderer.  A de-

tailed analysis of the hybrid lattice-reduction and meet-in-the-middle attack. Journal of Mathematical Cryptology, 13(1): 1–??, March 2019. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/
view/j/jmc.2019.13.issue-
1/jmc-2016-0044/jmc-2016-
0044.xml.

Yasuda:2020:SDD

[Yas20] Masaya Yasuda. Self-dual Deep-BKZ for finding short lattice vectors. Journal of Mathematical Cryptology, 14(1):84– 94, June 14, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

https://www.degruyter.com/

                                        document/doi/10.1515/jmc-
                                        2015-0053/html.

Yamashita:2024:BBI

[YH24] Kyosuke Yamashita and Keisuke Hara. On the black-box impossibility of multi-designated verifiers signature schemes from ring signature schemes. Journal of Mathematical Cryptology, 18(1):??, January 2024. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2023-0028/html.

Yoon:2015:NMC

[Yoo15] Kisoon Yoon. A new method of choosing primitive elements for Brezing–Weng families of pairing-friendly elliptic curves. Journal of Mathematical Cryptology, 9(1):1–9, 2015. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Yasuda:2014:EDR

[YYS^{+}14] Masaya Yasuda, Kazuhiro Yokoyama, Takeshi Shimoyama, Jun Kogure, and Takeshi Koshiba. On the exact decryption range for Gentry–Halevi’s implementation of fully homomorphic encryption. Journal of Mathematical Cryptology, 8(3): 305–329, 2014. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Yasuda:2017:ADS

[YYS^{+}17] Masaya Yasuda, Kazuhiro Yokoyama, Takeshi Shimoyama, Jun Kogure, and Takeshi Koshiba. Analysis of decreasing squared-sum of Gram–Schmidt lengths for short lattice vectors. Journal of Mathematical Cryptology, 11(1):1–24, 2017. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Yokoyama:2020:CBS

[YYTK20] Kazuhiro Yokoyama, Masaya Yasuda, Yasushi Takahashi, and Jun Kogure. Complexity bounds on Semaev’s naive index calculus method for ECDLP. Journal of Mathematical Cryptology, 14 (1):460–485, October 30, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.
com/document/doi/10.1515/
jmc-2019-0029/html.

Zajac:2013:NMS

[Zaj13] Pavol Zajac. A new method to solve MRHS equation systems and its connection to group factorization. Journal of Mathematical Cryptology, 7(4): 367–381, 2013. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

Zhou:2022:CCB

[ZHM^{+}22] Yu Zhou, Jianyong Hu, Xudong Miao, Yu Han, and Fuzhong Zhang. On the confusion coefficient of Boolean functions. Journal of Mathematical Cryptology, 16(1):1–13, January 2022. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

URL https://www.degruyter.

                                        com/document/doi/10.1515/
                                        jmc-2021-0012/html.

Zhang:2021:SBS

[ZLA21] Jing Zhang, Yuan Li, and John O. Adeyeye. Sensitivities and block sensitivities of elementary symmetric Boolean functions. Journal of Mathematical Cryptology, 15(1):434– 453, April 22, 2021. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/
                                        document/doi/10.1515/jmc-
                                        2020-0042/html.

Zhou:2020:CPB

[ZMD20] Yu Zhou, Daoguang Mu, and Xinfeng Dong. On cryptographic properties of (n + 1)-bit S-boxes constructed by known n-bit S-boxes. Journal of Mathematical Cryptology, 15(1):258– 265, December 8, 2020. CO-DEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic).

                                        URL https://www.degruyter.
                                        com/document/doi/10.1515/
                                        jmc-2020-0004/html.

Zhang:2020:PPV

[ZSN20] Liang Feng Zhang and Reihaneh Safavi-Naini. Privacypreserving verifiable delegation of polynomial and matrix functions. Journal of Mathematical Cryptology, 14(1):153– 171, July 3, 2020. CODEN ???? ISSN 1862-2976 (print), 1862-2984 (electronic). URL

                                        https://www.degruyter.com/

REFERENCES 50

document/doi/10.1515/jmc-
2018-0039/html.